Pink Floyd and solo^{∗}

Michael Patriksson^{†}

May 21, 1998

Contents

The Piper at the Gates of Dawn (5/8 1967)[6/–] 1 A Saucerful of Secrets (29/6 1968)[9/–] . . . . . . 2 More (27/7 1969)[9/153] . . . . . . . . . . . . . . 2 Ummagumma (25/10 1969)[5/74] . . . . . . . . . 2 Zabriskie Point (/1 1970) . . . . . . . . . . . . . 3 Atom Heart Mother (10/10 1970)[1/55] . . . . . 3 Relics (14/5 1971)[32/152] . . . . . . . . . . . . . 4 Meddle (11/11 1971)[3/70] . . . . . . . . . . . . . 4 Obscured by Clouds (3/6 1972)[6/46] . . . . . . . 4 The Dark Side of the Moon (24/3 1973)[2/1]{A

Piece for Assorted Lunatics; Eclipsed} … 5 Wish You Were Here (15/9 1975)[1/2] . . . . . . 5 Animals (23/1 1977)[2/2] . . . . . . . . . . . . . 5 The Wall (30/11 1979)[3/1] . . . . . . . . . . . . 6 Songs only in The Wall movie . . . . . . . . . . . 8 A Collection of Great Dance Songs (23/11

1981)[37/31] . . . . . . . . . . . . . . . . . . 8 the final cut (21/3 1983)[1/6] . . . . . . . . . . . 8 A Momentary Lapse of Rea-

son (8/9 1987)[3/3]{Of Promises Broken, Signs of Life, Delusions of Maturity} …. 9

Delicate Sound of Thunder (22/11 1988)[11/11] . 10 The Division Bell (30/3 1994)[1/1]{Awaken to

the Sense of Reality} . . . . . . . . . . . . . 10 Singles . . . . . . . . . . . . . . . . . . . . . . . . 11 Unreleased songs . . . . . . . . . . . . . . . . . . 11 The Pros and Cons of Hitchhiking (1984) …. 11 When the Wind Blows (/10 1986) . . . . . . . . 12 Radio KAOS (1987){Home} . . . . . . . . . . . . 13 Amused To Death (26/8 1992) . . . . . . . . . . 14 David Gilmour (25/5 1978) . . . . . . . . . . . . 15 About Face (1984) . . . . . . . . . . . . . . . . . 16

The Piper at the Gates of Dawn (5/8 1967)[6/–]

Astronomy Domine [Ba, Wr](Ba)

[Peter Jenner: ”Moon in both [houses]...Scorpio, [Arabian Skies], Libra...Pluto was not discovered until 1930...[two/ten] seconds to [ignition]...all systems satisfied...”]

Lime and limpid green, a second scene, A fight between the blue you once knew. Floating down the sound resounds Around the icy waters underground. Jupiter and Saturn Oberon Miranda and Titania Neptune Titan Stars can frighten...

[Peter Jenner: ”...just completed orbital...”] Blinding signs flap flicker flicker flicker Blam pow pow Stairway scare Dan dare

who’s there...

Lime and limpid green The sound surrrounds the icy waters under Lime and limpid green The sound surrrounds the icy waters Underground

Lucifer Sam [Ba](Ba){Percy the Ratcatcher}

Lucifer Sam Siam cat Always sitting by your side Always by your side That cat’s something I can’t explain

∗Each record is supplied with original release date and charting success in Britain/U.S. Every song is supplied with [·]: singer(s), in order of appearance, or indicating that the song is instrumental; (·): writer(s); {·}: working name(s) or other name(s) used in concerts.

†The research leading to this list was sponsored by nobody; this support is greatfully acknowledged.

Jennifer Gentle you’re a witch You’re the left side He’s the right side oh no That cat’s something I can’t explain

Lucifer go to sea Be a hip cat Be a ship’s cat Somewhere anywhere That cat’s something I can’t explain

Night prowling sifting sand Padding around on the ground He’ll be found When you’re around That cat’s something I can’t explain

Matilda Mother [Wr, Ba](Ba)

There was a king Who ruled the land His majesty Was in command With silver eyes the scarlet eagle Showered silver on the people Oh Mother tell me more

Why’d you have to leave me there Hanging in my infant air, waiting You only have to read the lines of Scribbly black and everything shines

Across the stream with wooden shoes Bells to tell the King the news A thousand misty riders Climb up higher once upon a time

Wondering and dreaming The words had different meaning Yes they did...

For all the time spent in that room The doll’s house darkness old perfume And fairy stories held me high On clouds of sunlight floating by Oh Mother tell me more Tell me more...

Flaming [Ba](Ba)

Alone in the Clouds all blue Lying on an eiderdown, yippee You can’t see me but I can you

Lazing in the foggy dew Sitting on a unicorn no fear You can’t hear me but I can you

Watching buttercups cup the light Sitting on a dandelion too much I won’t touch you but then I might

Swimming through the starlit skies Travelling by telephone Hey ho here we go Ever so high

Alone in the clouds all blue Lying on an eiderdown, yippee You can’t see me but I can you

Pow R. Toc H. [instr, Ba, Wr](Ba, Wa, Wr, Ma){The Pink Jungle (as part 4 of ’The Journey’)} Take Up Thy Stethoscope and Walk [Wa](Wa)

Doctor Doctor I’m in bed Aching head Doctor Doctor gold is lead Doctor Doctor choke on bread Doctor Doctor underfed Doctor Doctor gold is lead Doctor Doctor Jesus bled Doctor Doctor pain is red Doctor Doctor dark doom Gruel ghoul greasy spoon Used spoon June bloom

Music seems to help the pain Seems to motivate the brain Doctor kindly tell your wife t That I’m alive Flowers thrive Realise Realise Realise

Interstellar  Overdrive  [instr](Ba,

Wa, Wr, Ma){The Labyrinths of Auximines (middle section, as part 5 of ’The Journey’)} The Gnome [Ba](Ba)

I want to tell You a story ’Bout a little Man if I can A gnome named Grimble Gromble And little gnomes stay in their homes Eating, sleeping, drinking their wine

He wore a scarlet tunic A blue-green hood, it looked quite good He had a big adventure Amidst the grass, fresh air at last Wining, dining, biding his time...

And then one day Hooray, another way for gnomes to say Ooh my...

Look at the sky, look at the river Isn’t it good? Look at the sky, look at the river Isn’t it good?

Winding, finding places to go

And then one day Hooray, another way for gnomes to say Ooh my ooh my...

Chapter 24 [Ba](Ba)

All movement Is accomplished In six stages, And the seventh Brings return The seven is the number of the young light It forms when darkness is increased by one

Change return success Going and coming without error Action brings good fortune Sunset

The time is with the month of winter solstice When the change is due to come Thunder in the Earth, the course of Heaven Things cannot be destroyed once and for all

Change return success Going and coming without error Action brings good fortune Sunset Sunrise

All movement is accomplished in six stages And the seventh brings return The seven is the number of the young light It forms when darkness is increased by one

Change return success Going and coming without error Action brings good fortune Sunset Sunrise

The Scarecrow [Ba](Ba)

The black and green scarecrow As ev’ryone knows Stood with a bird on his hat And straw everywhere He didn’t care... He stood in a field where barley grows

His head did no thinking his arms didn’t move Except when the wind cut up rough And mice ran around on the ground He stood in a field where barley grows

The black and green scarecrow is sadder than me But now he’s resigned to his fate ’Cause life’s not unkind He doesn’t mind He stood in a field where barley grows

Bike [Ba](Ba)

I’ve got a bike, You can ride it if you like It’s got a basket, a bell that rings And things to make it look good I’d give it to you if I could, but I borrowed it

You’re the kind of girl that fits in with my world I’ll give you anything, everything if you want thing

I’ve got a cloak it’s a bit of a joke There’s a tear up the front it’s red and black I’ve had it for months If you think it could look good then I guess it should

You’re the kind of girl that fits in with my world I’ll give you anything, everything if you want thing

I know a mouse and he hasn’t got a house I don’t know why I call him Gerald He’s getting rather old but he’s a good mouse

You’re the kind of girl that fits in with my world I’ll give you anything, everything if you want thing

I’ve got a clan of gingerbread men Here a man, there a man, lots of gingerbread men Take a couple if you wish, they’re on the dish

You’re the kind of girl that fits in with my world I’ll give you anything, everything if you want thing

I know a room of musical tunes Some rhyme, some ching, most of them are clockwork Let’s go into the other room and make them work

Checked against audio and lyrics sheet (remastered Capitol CD edition) 950730

A Saucerful of Secrets (29/6 1968)[9/–]

Let There Be More Light [Gi, Wa](Wa)

Far far far away way People heard him say say I will find a way way There will come a day day Something will be done Then at last the mighty ship descending on a point of flame Made contact with the human race at Mildenhall

Now now now is the time time time To be be be aware Carter’s father saw it there And knew the Rhull revealed to him The living soul of Hereward the Wake

Oh my, something in my eye eye Something in the sky sky Waiting there for me The outer lock rolled slowly back The servicemen were heard to sigh For there revealed in flowing robes was Lucy in the sky

Oh oh did you ever? No no never ever will they I’ll say! Summoning his cosmic power And glowing slightly from his toes The psychic emanations flowed

Remember a Day [Wr](Wr)

Remember a day before today A day when you were young Free to play along with time Evening never comes

Sing a song that can’t be sung Without the morning’s kiss Queen you shall be if you wish Look for your king

Why can’t we play today? Why can’t we stay that way?

Climb your favourite apple tree Try to catch the sun Hide from your little brother’s gun Dream yourself away Why can’t we reach the sun? Why can’t we blow the years away? Blow away Blow away Remember Remember

Set the Controls for the Heart of the Sun [Wa](Wa)

Little by little the night turns around. Counting the leaves which tremble at dawn Lotuses lean on each other in yearning Over the hills a swallow is resting Set the controls for the heart of the sun

Over the mountain watching the watcher Breaking the darkness waking the grapevine Knowledge of love is knowledge of shadow Love is the shadow that ripens the wine Set the controls for the heart of the sun The heart of the sun, the heart of the sun

Witness the man who waves at the wall Making the shape of his question to heaven Whether the sun will fall in the evening Will he remember the lesson of giving? Set the controls for the heart of the sun The heart of the sun, the heart of the sun

Corporal Clegg [Gi, Wa, Ma](Wa)

Corporal Clegg had a wooden leg He won it in the war In nineteen forty four Corporal Clegg had a medal too In orange, red and blue He found it in the zoo Dear oh dear, are they really sad for me? Dear oh dear, will they really laugh

at me?

Missus Clegg, you must be proud of him Missus Clegg, another drop of gin? Corporal Clegg, umbrella in the rain He’s never been the same No one is to blame

Corporal Clegg received his medal in a dream From her Majesty, the Queen His boots were very clean Missus Clegg, you must be proud of him Missus Clegg, another drop of gin?

A Saucerful of Secrets, Part a: Something Else, b: Syncopated Pandemonium, c: Storm Signal, d: Celestial Voices [instr, Gi](Wa, Wr, Ma, Gi){The Massed Gadgets of Hercules; The End of the Beginning (d as part 7 of ’The Journey’)}

See Saw [Wr](Wr){The Most Boring Song I’ve Ever Heard Bar Two}

Marigolds are very much in love But he doesn’t mind Picking up his sister He makes his way in to the see-saw land All the way she smiles She goes up as he goes down down Sits on a stick in the river Laughter in his sleep Sister’s throwing stones hoping for a hit He doesn’t know, so then She goes up while he goes down, down

Another time, another day A brother’s way to leave Another time, another day Selling plastic flowers on a Sunday afternoon Picking up weeds She hasn’t got

the time to care All can see he’s not there She grows up for another man, and he’s down

Another time, another day A brother’s way to leave Another time, another day Another time, another day A brother’s way to leave

Jugband Blues [Ba](Ba)

It’s awfully considerate of you to think of me here And I’m almost obliged to you for making it clear that I’m not here And I never knew the moon could be so big And I never knew the moon could be so blue And I’m grateful that you threw away my old shoes And brought me here instead dressed in red

And I’m wondering who could be writing this song

I don’t care if the sun don’t shine And I don’t care if nothing is mine And I don’t care if I’m nervous with you I’ll do my loving in the Winter

And the sea isn’t green And I love the queen And what exactly is a dream? And what exactly is a joke?

Checked against lyrics sheet (Shine On box edition) 950826

More (27/7 1969)[9/153]

Cirrus Minor [Wa](Wa)

In a churchyard by a river Lazing in the haze of midday Laughing in the grasses And the graves Yellow bird you are not long in singing and in flying on In laughing and in leaving

Willow weeping in the water Waving to the river-daughters Swaying in the ripples and the reeds On a trip to Cirrus Minor Saw a crater in the Sun A thousand miles of moonlight later

The Nile Song [Gi](Wa)

I was standing by the Nile When I saw the lady smile I would take her out for a while For a while

Like tears that like a child How her golden hair was blowing wild Then she spread her wings to fly For to fly

Soaring high above the breezes Going always where she pleases She will make it to the island In the Sun I will follow in her shadow And I’ll watch her from my window One day I will catch her

She is calling from the deep Summoning my soul to endless sleep She is bound to drag me down Drag me down

Crying Song [Gi, Wr](Wa)

We smile and smile We smile and smile Laughter echoes in your eyes We climb and climb We climb and climb Foot falls softly in the pines

We cry and cry We cry and cry Sadness passes in a while We role and roll We role and roll Help me roll away the stone

Up the Khyber [instr](Ma, Wr)

Green is the Colour [Gi](Wa){The Beginning (as part 1 of ’The Journey’)}

Heavy hung the canopy of blue Shade my eyes and I can see you White is the light that shines Through the dress that you wore She lay in the shadow of a wave Hazy were the visions overplayed Sunlight on her eyes but moonshine Made her blind everytime Green is the colour of her kind Quickness of the eye Deceives the mind Envy is the bond between The hopeful and the damned

Cymbaline [Gi LP, Wa movie](Wa){Nightmare (as part 6 of ’The Man’)}

The path you tread is narrow And the drop is sheer and very high The ravens all are watching From a vantage point nearby Apprehension creeping Like a tube train up your spine Will the tightrope reach the end Will the final couplet rhyme And it’s high time Cymbaline It’s high time Cymbaline Please wake me

A butterfly with broken wings Is falling by your side The ravens all are closing in There’s nowhere you can hide Your manager and agent Are both busy on the phone Selling coloured photographs To magazines back home And it’s high time Cymbaline It’s high time Cymbaline Please wake me

The lines converging where you stand They must have moved the picture plain The leaves are heavy round your feet You hear the thunder of the train Suddenly it strikes you That they’re moving into range And Doctor Strange Is always changing size And it’s high time Cymbaline It’s high time Cymbaline Please wake me And it’s high time Cymbaline It’s high time Cymbaline Please wake me

Party Sequence [instr](Wa, Wr, Gi, Ma)

Main Theme [instr](Wa, Wr, Gi, Ma)

Ibiza Bar [Gi](Wa, Wr, Gi, Ma)

I’m so afraid Of mistakes that I’ve made Shaking every time that I awake I feel like a cardboard cut-out man So build me a time When the characters rhyme And the story line is kind

I’ve aged and aged since the first page I’ve lived every line that you wrote Take me down take me down From the shelf above your head And build me a time When the characters rhyme And the story line is kind

I live where I’m left On the shelf like the rest And the epilogue reads like a sad song Please pick up your camera And use me again and build me a time When the characters rhyme And the story line is kind

Yea

More Blues [instr](Wa, Wr, Gi, Ma)

Quicksilver [instr](Wa, Wr, Gi, Ma){Sleeping (as part 5 of ’The Man’)}

A Spanish Piece [Gi](Gi)

Pass the tequila, Manuel (glug, snort) Listen, gringo, laugh at my lisp and I kill you I think This Spanish music It sets my soul on fire Lovely seniorita Your eyes are like stars And your teeth are like pearls Your ruby lips seniorita

Dramatic Theme [instr](Wa, Wr, Gi, Ma)

Checked against audio and lyrics sheet (remastered EMI CD edition) 960309

Ummagumma (25/10 1969)[5/74]

Astronomy Domine [Gi, Wr, Wa](Ba)

Lime and limpid green a second scene a fight between the blue you once knew Floating down the sound resounds around the icy waters underground Jupiter and Saturn Oberon, Miranda and Titania, Neptune Titan stars can frighten

Blinding signs flat flicka flicka flicka blam pow Stairway, scale down, dare whose lair

Lime and limpid green, the sound surrounds the icy waters underground Lime and limpid green, the sound surrounds the icy waters underground

Careful With That Axe, Eugene [Wa, Gi](Wa, Wr, Ma, Gi){Keep Smiling People; Beset by Creatures of the Deep (as part 2 of ’The Journey); Come in Number 51, Your Time Is Up}

Careful with that axe, Eugene Ahhhhhhhhhhhhh (repeat)

Set the Controls for the Heart of the Sun [Wa](Wa)

Little by little the night turns around Counting the leaves which tremble at dawn Novices lean on each other in yearning Under the leaves the swallow is resting

Set the controls for the heart of the sun

Over the mountain watching the watcher Breaking the darkness waking the grapevine One inch of love is one inch of shadow Love is the shadow that ripen’s the wine

Set the controls for the heart of the sun The heart of the sun(repeat)

Witness the man who raves at the wall Making the shape of his question to heaven Whether the sun will fall in the evening Will he remember the lesson of giving

Set the controls for the heart of the sun The heart of the sun(repeat)

A Saucerful of Secrets, Part a: Something Else, b: Syncopated Pandemonium, c: Storm Signal, d: Celestial Voices [instr](Wa, Wr, Ma, Gi){The Massed Gadgets of Hercules; The End of the Beginning (d as part 7 of ’The Journey’)}

Sysyphus, Parts 1–4 [Parts a–c instr, Wr](Wr)

Grantchester Meadows [Wa](Wa){Daybreak (as part 1 of ’The Man’)}

Icy wind of night, be gome. This is not your domain. In the sky a bird was heard to cry. Misty morning whisperings and gentle stirring sounds Belied a deathly silence that lay all around. Hear the lark and harken to the barking of the dog fox gone to ground. See the splashing of the kingfisher flashing to the water. And a river of green is sliding unseen beneath the trees, Laughing as it passes through the endless summer making for the sea.

In the lazy water meadow I lay me down. All around me, Golden sunflakes settle on the ground, Basking in the sunshine of a by gone afternoon, Bringing sounds of yesterday into this city room. Hear the lark and harken to the barking of the dog fox gone to ground. See the splashing of the kingfisher flashing to the water. And a river of green is sliding unseen beneath the trees, Laughing as it passes through the endless summer making for the sea.

In the lazy water meadow I lay me down. All around me, Golden sunflakes covering the ground, Basking in the sunshine of a by gone afternoon, Bringing sounds of yesterday into my city room. Hear the lark and harken to the barking of the dog fox gone to ground. See the splashing of the kingfisher flashing to the water. And a river of green is sliding unseen beneath the trees, Laughing as it passes through the endless summer making for the sea.

Several Species of Small Furry Animals Gathered Together in a Cave and Grooving with a Pict [instr, Wa](Wa)

[Wa with a Scottish-like accent: ”Aye an’ a bit of Mackeral settler rack and ruin ran it doon by the haim, ’ma place well I slapped me and I slapped it doon in the side and I cried, cried, cried.

The fear a fallen down taken never back the raize and then Craig Marion, get out wi’ ye Claymore out mi pocket a’ ran doon, doon the middin stain picking the fiery horde that was fallen around ma feet.

’Never,’ he cried, ’never shall it ye get me alive ye rotten hound of the burnie crew.’ Well I snatched fer the blade O my Claymore cut and thrust and I fell doon before him round his feet. Aye!

A roar he cried frae the bottom of his heart that I would nay fall but as dead, dead as ’a can be by his feet; de ya ken?

...and the wind cried back.”]

[Wa: ”Thank you” (left channel)]

(Rumour has it [Amazing Pudding] that various oddities can be heard by playing the song [vinyl version!] at different speeds. [I will check this out eventually.] This includes:

Wa: ”Bring back my guitar” (repeated) [78rpm]

Gi: ”That was pretty avant-garde, wasn’t it?” [16rpm])

The Narrow Way, Parts 1–3 [Parts 1 and 2 instr, Gi](Gi){ (part 3 of ’The Journey’) Baby Blue Shuffle in D Major}

Following the path as it leads towards the Darkness in the North Passing strangers’ faces show their sympathy They’ve seen that hope before And if you want to stay for a little bit Rest your aching limbs for a little bit Before you the night is beckoning And you know you can’t delay You hear the night birds calling you But you can’t catch the words they say Close your aching eyes (?) beyond your name

Mist is swelling, creatures crawling, Hear the roar grow louder in your ears You know the folly was your own But the force behind can’t conquer all your fears And if you want to stay for a little bit Rest your aching limbs for a little bit Before you the night is beckoning And you know you can’t delay You hear the night birds calling you But you can’t catch the words they say Close your aching eyes (?) beyond your name

Cast your thoughts back many years To the time when there was light in every dawning Perhaps a time will come When the nights will be as bright as on that morning And if you want to stay for a little bit Rest your aching limbs for a little bit Before you the night is beckoning And you know you can’t delay You hear the night birds calling you But you can’t catch the words they say Close your aching eyes (?) beyond your name Rest your weary head, be on your way.

The Vizier’s Garden Party, Part 1: Entrance, Part 2: Entertainment, Part 3: Exit [instr](Ma)

Zabriskie Point (/1 1970)

Heart Beat, Pig Meat [instr](Wa, Wr, Gi, Ma)

Crumbling Land [Gi](Wa, Wr, Gi, Ma)

In a while I’ll find the time to make the sunshine mine In a smile I saw a single needle in the sky Wheeling, soaring, gliding by

On a hill a little man with many shining things Shiny pool, a shiny car and shiny diamond rings Wining, dining, shining king

And the eagle flies in clear blue skys Breathing in the clear blue air Back here on the ground another dealer coughs and dies And fifty more come rolling off The Ford production line

Then a man appearing like up in our on the sand In his hand a moving picture of the crumbling land Screaming, scheming, movie man

Here we go, hold your nose and see if something blows Close your eyes, count to ten and see the sunrise rise Climbing, higher, to the sky

Climbing all the birds a lake the dove the lies On the finger of the king On high the eagle spies the glitter of a gun And wheeling in a climbing turn he flies into the sun

Come in Number 51, Your Time Is Up [instr, Wa](Wa, Wr, Gi, Ma){Careful with That Axe, Eugene; Beset by Creatures of the Deep (as part 2 of ’The Journey)}

Atom Heart Mother (10/10 1970)[1/55]

Atom Heart Mother, Part a: Father’s Shout, Part b: Breast Milky, Part c: Mother Fore, Part d: Funky Dung, Part e: Mind Your Throats Please, Part f: Remergence [instr, feat. the John Aldiss Choir](Ma, Gi, Wa, Wr, Geesin){The Amazing Pudding}

[(at 17:28:) ”Here is a loud announcement” (at 19:08:) ”Silence in the studio!”]

If [Wa](Wa)

If I were a swan, I’d be gone. If I were a train, I’d be late. And if I were a good man, I’d talk with you more often than I do. If I were to sleep, I could dream. If I were afraid, I could hide. If I go insane, please don’t put your wires in my brain.

If I were the moon, I’d be cool. If I were a road, I would bend. If I were a good man, I’d understand the spaces between friends. If I were alone, I would cry. And if I were with you, I’d be home and dry. And if I go insane, Will you still let me join in with the game?

If I were a swan, I’d be gone. If I were a train, I’d be late again. If I were a good man, I’d talk to you more often than I do.

Summer ’68 [Wr](Wr)

Would you like to something before you leave? Perhaps you’d care to state exactly how you feel. We say goodbye before we’ve said hello. I hardly even like you. I shouldn’t care at all. We met just six hours ago. The music was too loud. From your bed I came today and lost a bloody year. And I would like to know, how do you feel? How do you feel?

Not a single word was said. They lied still without fears. Occasionally you showed a smile, but what was the need? I felt the cold far too soon in a wind of ninetyfive. My friends are lying in the sun, I wish I was there. Tomorrow brings another town, another girl like you. Have you time before you leave to greet another man Just to let me know, how do you feel? How do you feel?

Goodbye to you. Childish bangles too. I’ve had enough for one day.

Fat Old Sun [Gi](Gi)

When the fat old sun in the sky is falling, Summer evening birds are calling. Summer’s thunder time of year, The sound of music in my ears. Distant bells, New mown grass smells so sweet. By the river holding hands, Roll me up and lay me down. And if you sit, Don’t make a sound. Pick your feet up off the ground. And if you hear as the warm night falls The silver sound from a time so strange, Sing to me, sing to me.

When that fat old sun in the sky is falling, Summer evening birds are calling. Children’s laughter in my ears, The last sunlight disappears.

And if you sit, Don’t make a sound. Pick your feet up off the ground. And if you hear as the warm night falls The silver sound from a time so strange, Sing to me, sing to me.

Alan’s Psychedelic Breakfast, Part a: Rise and Shine, Part b: Sunny Side Up, Part c: Morning Glory [instr, feat. Alan Stiles] (Wa, Ma, Gi, Wr)

Relics (14/5 1971)[32/152]

Arnold Layne [Ba](Ba)

Interstellar  Overdrive  [instr](Ba,

Wa, Wr, Ma){The Labyrinths of Auximines (middle section, as part 5 of ’The Journey’)

See Emily Play [Ba](Ba){Games for May}

Remember a Day [Wr](Wr)

Paint Box [Wr](Wr)

Julia Dream [Wa](Wa){Doreen’s Dream}

Careful With That Axe, Eugene [instr, Wa] (Wa, Wr, Ma, Gi){Keep Smiling People; Beset by Creatures of the Deep (as part 2 of ’The Journey); Come in Number 51, Your Time Is Up}

Cirrus Minor [Wa](Wa)

The Nile Song [Gi](Wa)

Biding My Time [Wa](Wa){Work (as part 2 of ’The Man’)}

Wasting my time, Resting my mind And I’ll never pine For the sad days and the bad days When we was workin’ from nine to five And if you don’t mind I’ll spend my time Here by the fire side In the warm light of her eyes And if you don’t mind I’ll spend my time Here by the fire side In the warm light of her eyes

Bike [Ba](Ba)

Meddle (11/11 1971)[3/70]

One of These Days [Ma](Wa, Wr, Gi, Ma)

One of these days I’m going to cut you into little pieces

A Pillow of Winds [Gi](Wa, Gi)

A cloud of eiderdown draws around me Softening the sound Sleepy time when I lie with my love by my side And she’s breathing low, and the candle dies

When night comes down you lock the door The book falls to the floor As darkness falls and waves roll by The seasons change the wind is warm

Now wakes the owl, now sleeps the swan Behold a dream, the dream is gone Green fields, a cold rain Is falling in a golden dawn

And deep beneath the ground the early morning sounds And I go down Sleepy time when I lie with my love by my side And she’s breathing low

And I rise like a bird In the haze when the first rays touch the sky And the night winds die

Fearless [Gi](Wa, Gi)

You say the hill’s too steep to climb, climbing You say you’d like to see me try, climbing You pick the place and I’ll choose the time And I’ll climb the hill in my own way Just wait a while for the right day And as I rise above the tree-line and the clouds I look down Hearing the sound Of the things you’ve said today

Fearlessly the idiot faced the crowd, smiling Merciless the magistrate turns round, frowning And who’s the fool who wears the crown No doubt in your own way And every day is the right day And as you rise above the fear-lines in his frown You look down hear the sound of the faces in the crowd

[Interposing “You’ll Never Walk Alone” (Rodgers and Hammerstein), sung by Liverpool fans at the Kop end of Liverpool F.C.’s Anfield ground]

San Tropez [Wa](Wa)

As I reach for a peach Slide a line down behind a sofa in San Tropez Breaking a stick with a brick on the sand Riding a wave in the wake of an old sedan Sleeping alone in the drone of the darkness Scratched by the sand that fell from my love Deep in my dreams and I still hear her calling If you’re alone, I’ll come ho-ho-home

Backward and homebound, the pidgeon, the dove Gone with the wind and the rain on an airplane Born in a home with no silver spoon I’m drinking champagne like a big tycoon Sooner than wait for a break in the weather I’ll gather my far flung thoughts together Speeding away on a wind to a new day If your alone, I’ll come ho-ho-home

And I pause for a while by a country stile And listen to the things they say Digging for gold in a hole in my hand Open the book, take a look at the way things stand And you’re leading me down to the place by the sea I hear your soft voice calling to me Making a date for later by phone And if you’re alone I’ll come home

Seamus [Gi](Wa, Wr, Gi, Ma){Mademoiselle Nobs (at the Pompeii performance)}

I was in the kitchen Seamus, that’s the dog, was outside Well I was in the kitchen Seamus, my old hound, was outside Well you know the sun was sinking slowly And my old hound-dog sat right down and cried

[(According to rumour (still not checked): ”Here is the real dog” is heard just before one of the dog’s ’solos’]

Echoes [Gi, Wr](Wa, Wr, Gi, Ma){Nothing, Parts 1–24; The Son of Nothing; The Return of the Son of Nothing; Nothing To The Sun And Back (working titles); We Won the Double; Through The Knotholes In Granny’s Wooden Leg; The March of the Dambusters (as introduced by Wa during concerts)}

Overhead the albatross hangs motionless upon the air And deep beneath the rolling waves In labyrinths of coral caves The echo of a distant tide Comes willowing across the sand And everything is green and submarine

And no-one showed us to the land And no-one knows the wheres or whys But something stirs and Something tries And starts to climb towards the light

Strangers passing in the street By chance two separate glances meet And I am you and what I see is me And do I take you by the hand And lead you through the land And help me understand the best I can

And no-one calls us to move on And no-one forces down our eyes And no-one speaks And no-one tries And no-one flies around the sun

Cloudless every day you fall upon my waking eyes Inviting and inciting me to rise And through the window in the wall Come streaming in on sunlight wings A million bright ambassadors of morning

And no-one sings me lullabies And no-one makes me close my eyes So I throw the windows wide And call to you across the sky

Checked against audio (MFSL CD edition) and lyrics sheet (Shine On box edition) 950825

Obscured by Clouds (3/6 1972)[6/46]

Obscured by Clouds [instr](Wa, Gi) When You’re In [instr](Wa, Gi, Wr, Ma) Burning Bridges [Gi, Wr](Wa, Wr)

Bridges burning gladly Merging with the shadows Flickering between the lines Stolen moments Floating softly on the air Borne on wings of fire And climbing higher Ancient bonds are breaking Moving on and changing sides Dreaming of a new day Cast aside the other way Magic visions stirring Kindled by and burning flames Rise in her eyes

The door stands ajar The walls that once were high Beyond the gilded cage Beyond the reach of time The moment is at hand She breaks the golden band

The Gold it’s in the... [Gi, Wr](Wa, Gi)

Come on my friends let’s make for the hills They say there’s gold and I’m looking for thrills You can get your hands on whatever we find ’Cause I’m only coming along for the ride

Well you go on your way and I’ll go mine I don’t care if we get there on time Everybody’s searching for something they say I’ll get my kicks on the way

Over the mountains across the seas Who knows what may be waiting for me I could sail forever to strange sounding names Faces of people and places don’t change

All I have to do is just close my eye to see The sea gulls wheeling in the far distant skies All I wanna tell you all I wanna say Is count me in on the journey Don’t expect me to stay

Wot’s ... uh the Deal [Gi](Wa, Gi)

Heaven sent the promised land Looks alright from where I stand ’Cause I’m the man On the outside looking in Waiting on the first step Show me where the key is kept Point me down the right line Because it’s time

To let me in from the cold Turn my lead into gold ’Cause there’s a chill wind Blowing in my soul And I think I’m growing old

Flash ... the readies ... wot’s ... uh the deal Got to make it to the next meal Try to keep up with the turning of the wheel Mile after mile stone after stone You turn to speak but you’re alone Million miles from home You’re on your own

So let me in from the cold Turn my lead into gold ’Cause there’s a chill wind Blowing in my soul And I think I’m growing old

Fire bright by candle light With her by my side If she prefers we need never stir again Someone sent the promised land Well I grabbed it with both hands Now I’m the man On the inside looking out

Hear me shout come on in What’s the news where you’ve been ’Cause there’s no wind left in my soul And I’ve grown old

Mudmen [instr](Wr, Gi) Childhood’s End [Gi](Gi)

You shout in your sleep Perhaps the price was just too steep Is your conscience at rest if once put to the test You awake with a start To just the beating of your heart Just one man beneath the sky Just two ears just two eyes

You set sail across the sea Of long past thoughts and memories Childhood’s end your fantasies Merge with harsh realities And then as the sail is hoist You find your eyes are growing moist And all the fears never voiced say you have to make the final choice

Who are you and who am I To say we know the reason why Some are born some men die Beneath one infinite sky There’ll be war and there’ll be peace But everything one day will cease All the iron turned to rust All the proud men turned to dust So all things time will mend So this song will end

Free Four [Wa](Wa)

One two three four

The memories of a man in his old age Are the deeds of a man in his prime You shuffle in the gloom of the sick room And talk to yourself as you die

And life is a short warm moment And death is a long cold rest You get your chance to try In the twinkling of an eye Eighty years with luck or even less

So all aboard for the American tour And maybe you’ll make it to the top But mind how you go And I can tell you ’cause I know You may find it hard to get off

But you are the angel of death And I am the dead man’s son He was buried like a mole in a fox-hole And everyone’s still on the run And who is the master of foxhounds And who says the hunt has begun And who calls the tune in the courtroom And who beats the funeral drum

The memories of a man in his old age Are the deeds of a man in his prime You shuffle in the gloom of the sick room And talk to yourself as you die

Stay [Wr, Wa](Wr, Wa)

Stay And help me to end the day And of you don’t mind We’ll break a bottle of wine Stick around And maybe we’ll put one down ’Cause I wanna find What lies behind those eyes Midnight blue burning gold A yellow moon is growing cold

I rise Looking through my morning eyes Surprised to find You by my side Rack my brain and Try to remember your name To find the words To tell you goodbye Morning dues new born day Midnight blue turn to grey

Midnight blue burning gold A yellow moon is growing cold

Absolutely Curtains [instr](Wa, Gi, Wr, Ma)

[From the film: New Guinea natives singing]

Checked against audio and lyrics sheet (remastered EMI CD edition) 960309

The Dark Side of the Moon (24/3 1973)[2/1]{A Piece for Assorted Lunatics; Eclipsed}

Speak to Me [instr](Ma)

[Roger the Hat: I’ve been mad for fuckin’ years, absolutely years ... been over the edge for yonks, been working me buns off for bands...

Man: ”I’ve always been mad, I know I’ve been mad like the most of us have. They have to explain why you’re mad, even if you’re not mad.”]

Breathe in the Air [Gi](Wa, Gi, Wr)

Breathe, breathe in the air Don’t be afraid to care Leave but don’t leave me Look around and choose your own ground For long you live and high you fly And smiles you’ll give and tears you’ll cry And all you touch and all you see Is all your life will ever be

Run, run rabbit run Dig that hole, forget the sun, And when at last the work is done Don’t sit down it’s time to dig another one For long you live and high you fly But only if you ride the tide And balanced on the biggest wave You race towards an early grave

On the Run [instr](Gi, Wa)

[Female announcer, announcing flights at airport, including ’Rome’] [”Live for today, gone tomorrow, that’s me, HaHaHaaaaaa!”]

Time [Gi, Wr](Ma, Wa, Wr, Gi)

Ticking away the moments that make up a dull day You fritter and waste the hours in an off hand way Kicking around on a piece of ground in your home town Waiting for someone or something to show you the way

Tired of lying in the sunshine staying home to watch the rain You are young and life is long and there is time to kill today And then the one day you find ten years have got behind you No one told you when to run, you missed the starting gun

And you run and you run to catch up with the sun, but it’s sinking And racing around to come up behind you again The sun is the same in the relative way, but you’re older Shorter of breath and one day closer to death

Every year is getting shorter, never seem to find the time Plans that either come to naught or half a page of scribbled lines Hanging on in quiet desparation is the English way The time is gone the song is over, thought I’d something more to say

Breathe Reprise [Gi](Ma, Wa, Wr, Gi)

Home, home again I like to be here when I can When I come home cold and tired It’s good to warm my bones beside the fire Far away across the field The tolling of the iron bell Calls the faithful to their knees To hear the softly spoken magic spells

The Great Gig in the Sky [Clare Torry](Wr){The Mortality Sequence; Ecclesiastics}

[Man: ”And I am not frightened of dying; any time will do, I don’t mind. Why should I be frightened of dying? There’s no reason for it; you got to go some time.”]

[Man: ”If you can hear this whispering you are dying.”] [Man: ”I never said I was frightened of dying.”]

Money [Gi](Wa)

Money, get away Get a good job with more pay and your O.K. Money it’s a gas Grab that cash with both hands and make a stash New car, caviar, four star daydream, Think I’ll buy me a football team

Money get back I’m all right Jack keep your hands off my stack. Money it’s a hit Don’t give me that do goody good bullshit I’m in the hi-fidelity first class travelling set And I think I need a Lear jet

Money it’s a crime Share it fairly but don’t take a slice of my pie Money so they say Is the root of all evil today But if you ask for a rise it’s no surprise that they’re giving none away

[Man: ”He he, I was in the right.” Man: ”Yes, absolutely in the right.” Man: ”I certainly was in the right.” Woman: ”You was definitely in the right. That geezer was cruising for a bruis-

ing!...I guess he was cruisin’ for bruisin’.”

Man: ”Why does anyone do anything?” Man: ”Yeah!” Man: ”I don’t know, I was really drunk at the time.” Man: ”I was just telling him, he couldn’t get into number 2. He was asking why

he wasn’t coming up on freely, after I was yelling and screaming and telling him why he wasn’t coming up on freely. It came as a heavy blow, but we sorted the matter out.”]

Us and Them [Wa, Gi, Wr](Wa, Wr){The Violent Sequence}

Us, and them And after all we’re only ordinary men Me, and you God only knows it’s not what we would choose to do Forward he cried from the rear and the front rank died And the General sat, and the lines on the map moved from side to side

Black and blue And who knows which is which and who is who Up and Down And in the end it’s only round and round and round Haven’t you heard it’s a battle of words the poster bearer cried Listen son, said the man with the gun There’s room for you inside

[Roger the Hat: ”I mean, they’re not gunna kill ya, so if you give ’em a quick short, sharp, shock, they won’t do it again. Dig it? I mean he get off lightly, ’cos I would’ve given him a thrashing - I only hit him once! It was only a difference of opinion, but really ... I mean good manners don’t cost nothing do they, eh?”]

Down and Out It can’t be helped but there’s a lot of it about With, without And who’ll deny it’s what the fighting’s all about Get out of the way, it’s a busy day And I’ve got things on my mind For want of the price of tea and a slice The old man died

Any Colour You Like [instr](Gi, Ma, Wr) Brain Damage [Wa](Wa){The Dark Side of the Moon}

The lunatic is on the grass The lunatic is on the grass Remembering games and daisy chains and laughs Got to keep the loonies on the path

The lunatic is in the hall The lunatics are in my hall The paper holds their folded faces to the floor And every day the paper boy brings more

And if the dam breaks open many years too soon And if there is no room upon the hill And if your head explodes with dark forbodings too I’ll see you on the dark side of the moon

The lunatic is in my head The lunatic is in my head You raise the blade, you make the change You re-arrange me ’till I’m sane

You lock the door And throw away the key There’s someone in my head but it’s not me.

And if the cloud bursts, thunder in your ear You shout and no one seems to hear And if the band you’re in starts playing different tunes I’ll see you on the dark side of the moon

[Man: ”I can’t think of anything to say except... I think it’s marvelous! Ha-HaHa!”]

Eclipse [Wa, Gi](Wa)

All that you touch And all that you see All that you taste All you feel And all that you love And all that you hate All you distrust All you save And all that you give And all that you deal And all that you buy beg, borrow or steal And all you create And all you destroy And all that you do And all that you say And all that you eat and everyone you meet And all that you slight and everyone you fight And all that is now And all that is gone And all that’s to come and everything under the sun is in tune but the sun is eclipsed by the moon

[Jerry Driscoll: ”There’s no dark side in the moon, really. Matter of fact, it’s all dark.”]

[Very faint: string instruments]

Checked against audio and lyrics sheet (MFSL CD edition) 950618

Wish You Were Here (15/9 1975)[1/2]

Shine on You Crazy Diamond, Parts I–V [Wa](Wr, Wa, Gi; Gi, Wa, Wr; Wa, Gi, Wr; Gi, Wr, Wa; Wa)

Remember when you were young, you shone like the sun. Shine on you crazy diamond. Now there’s a look in your eyes, like black holes in the sky. Shine on you crazy diamond. You were caught in the crossfire of childhood and stardom, blown on the steel breeze. Come on you target for faraway laughter, come on you stranger, you legend, you martyr, and shine!

You reached for the secret too soon, you cried for the moon. Shine on you crazy diamond. Threatened by shadows at night, and exposed in the light. Shine on you crazy diamond. Well you wore out your welcome with random precision, rode on the steel breeze. Come on you raver, you seer of visions, come on you painter, you piper, you prisoner, and shine!

Welcome to the Machine [Wa, Gi](Wa)

Welcome my son, welcome to the machine. Where have you been? It’s alright we know where you’ve been. You’ve been in the pipeline, filling in time, provided with toys and ’Scouting for Boys’. You bought a guitar to punish your ma, you didn’t like school, and you know you’re nobody’s fool, So welcome to the machine.

Welcome my son, welcome to the machine. What did you dream? It’s alright we told you what to dream. You dreamed of a big star, he played a mean guitar, He always ate at the Steak Bar. He loved to drive in his Jaguar. So welcome to the machine.

Have a Cigar [Roy Harper](Wa)

Come in here, dear boy, have a cigar. You’re gonna go far, you’re gonna fly high, you’re never gonna die, you’re gonna make it if you try; they’re gonna love you. Well I’ve always had a deep respect, and I mean that most sincerely. The band is just fantastic, that is really what I think. Oh by the way which one’s Pink? And did we tell you the name of the game, boy, we call it Riding the Gravy Train.

We’re just knocked out. We heard about the sell out. You gotta get an album out, You owe it to the people. We’re so happy we can hardly count. Everybody else is just green, have you seen the chart? It’s a helluva start, it could be made into a monster if we all pull together as a team. And did we tell you the name of the game, boy, we call it Riding the Gravy Train.

Wish You Were Here [Gi](Wa, Gi)

So, so you think you can tell Heaven from Hell, blue skies from pain. Can you tell a green field from a cold steel rail? A smile from a veil? Do you think you can tell?

And did they get you to trade your heroes for ghosts? Hot ashes for trees? Hot air for a cool breeze? Cold comfort for change? And did you exchange a walk-on part in the war for a lead role in a cage?

How I wish, how I wish you were here. We’re just two lost souls swimming in a fish bowl, year after year, Running over the same old ground. What have you found? The same old fears. Wish you were here.

Shine on You Crazy Diamond, Parts VI–IX [Wa](Wr, Wa, Gi; Wa, Gi, Wr; Gi, Wa, Wr; Wr)

Nobody knows where you are, how near or how far. Shine on you crazy diamond. Pile on many more layers and I’ll be joining you there. Shine on you crazy diamond. And we’ll bask in the shadow of yesterday’s triumph, sail on the steel breeze. Come on you boy-child, you winner and loser, come on you miner for truth and delusion, and shine!

Checked against audio and lyrics sheet (SONY SBM CD edition) 950825

Animals (23/1 1977)[2/2]

Pigs on the Wing, Part 1 [Wa](Wa)

If you didn’t care what happened to me And I didn’t care for you We would zig zag our way through the boredom and pain Occasionally glancing up through the rain Wondering which of the buggers to blame And watching for pigs on the wing.

Dogs [Wa, Gi](Gi, Wa){Gotta Be Crazy}

You gotta be crazy, you gotta have a real need You gotta sleep on your toes, and when you’re on the street You gotta be able to pick out the easy meat with your eyes closed And then moving in silently, down wind and out of sight You gotta strike when the moment is right without thinking

And after a while, you can work on points for style Like the club tie, and the firm handshake A certain look in the eye and an easy smile You have to be trusted by the people that you lie to So that when they turn their backs on you You’ll get the chance to put the knife in

You gotta keep one eye looking over your shoulder You know it’s going to get harder and harder and harder as you get older Yeah And in the end you’ll pack up, and fly down south Hide your head in the sand Just another sad old man All alone and dying of cancer

And when you lose control, you’ll reap the harvest You have sown And as the fear grows, the bad blood slows and turns to stone And it’s too late to loose the weight you used to need to throw around So have a good drown, as you go down, all alone Dragged down by the stone

I gotta admit that I’m a little bit confused Sometimes it seems to me as if I’m just being used Gotta stay awake, gotta try and shake off this creeping malaise If I don’t stand my own ground, how can I find my way out of this maze?

Deaf dumb and blind, you just keep on pretending That everyone’s expendable and no one has a real friend And it seems to you the thing to do would be to isolate the winner And everything’s done under the sun, and you believe at heart, everyone’s a killer

Who was born in a house full of pain Who was trained not to spit in the fan Who was told what to do by the man Who was broken by trained personnel Who was fitted with collar and chain Who was given a pat on the back Who was breaking away from the pack Who was only a stranger at home Who was ground down in the end Who was found dead on the phone Who was dragged down by the stone Who was dragged down by the stone

Pigs (Three different ones) [Wa](Wa)

Big man, pig man, ha ha charade you are Wwhuuuu! You well heeled big wheel, ha ha charade you are And when your hand is on your heart You’re nearly a good laugh Almost a joker With your head down in the pig bin Saying ”Keep on digging” Pig stain on your fat chin What do you hope to find? Down in the pig mine You’re nearly a laugh You’re nearly a laugh But you’re really a cry

Bus stop rat bag, ha ha charade you are You fucked up old hag, ha ha charade you are You radiate cold shafts of broken glass You’re nearly a good laugh Almost worth a quick grin You like the feel of steel You’re hot stuff with a hat pin And good fun with a hand gun You’re nearly a laugh You’re nearly a laugh But you’re really a cry

Hey you, Whitehouse, ha ha charade you are You house proud town mouse, ha ha charade you are You’re trying to keep our feelings off the street You’re nearly a real treat All tight lips and cold feet And do you feel abused? .....! .....! .....! .....! You gotta stem the evil tide And keep it all on the inside Mary, you’re nearly a treat Mary, you’re nearly a treat But you’re really a cry

Sheep [Wa](Wa){Raving and Drooling}

Harmlessly passing your time in the grassland away Only dimly aware of a certain unease in the air You better watch out There may be dogs about I’ve looked over Jordan and I have seen Things are not what they seem

What do you get for pretending the danger’s not real Meek and obedient you follow the leader Down well-trodden corridors into the valley of steel What a surprise! A look of terminal shock in your eyes Now things are really what they seem No, this is no bad dream

The Lord is my shepherd, I shall not want He makes me down to lie Through pastures green he leadeth me the silent waters by With bright knives he releaseth my soul He maketh me to hang on hooks in high places He converteth me to lamb cutlets For lo, he hath great power, and great hunger When cometh the day we lowly ones Through quiet reflection and great dedication Master the art of karate Lo, we shall rise up And then we’ll make the bugger’s eyes water

Bleating and babbling we fell on his neck with a scream Wave upon wave of demented avengers March cheerfully out of obscurity into the dream

Have you heard the news? The dogs are dead! You better stay home And do as you’re told Get out of the road if you want to grow old

Pigs on the Wing, Part 2 [Wa](Wa)

You know that I care what happens to you And I know that you care for me too So I don’t feel alone Or the weight of the stone Now that I’ve found somewhere safe To bury my bone And any fool knows a dog needs a home A shelter from pigs on the wing

Checked against audio and lyrics sheet (Shine On box edition) 950825

The Wall (30/11 1979)[3/1]

In the Flesh? [Wa](Wa)

[Man: ”...we came in?”] So ya Thought ya Might like to go to the show To feel the warm thrill of confusion That space cadet glow Tell me is something eluding you sunshine? Is this not what you expected to see? If you wanna find out what’s behind these cold eyes You’ll just have to claw your way through this Disguise

[Man: ”Lights! Roll the sound effects! Action! Drop it! Drop it on ’em! Drop it on them!!!!!”]

The Thin Ice [Gi, Wa](Wa)

Mamma loves her baby And daddy loves you too And the sea may look warm to you babe, And the sky may look blue Ooooh babe Oooh baby blue Ooooooh Babe

If you should go skating On the thin ice of modern life Draggin’ behind you the silent reproach Of a million tear stained eyes Don’t be surprised when a crack in the ice Appears under your feet You slip out of your depth and out of your mind With your fear flowing out behind you As you claw the thin ice

Another Brick in the Wall, Part 1 [Wa](Wa)

Daddy’s flown across the ocean Leaving just a memory A snapshot in the family album Daddy what else did you leave for me? Daddy what d’ya leave behind for me?

All in all it was just a brick in the wall All in all it was all just bricks in the wall

[Teacher: ”You! Yes, you! Stand still laddie!”]

The Happiest Days of Our Lives [Wa](Wa)

When we grew up and went to school There were certain teachers who would Hurt the children in any way they could

[”OOF!” (someone being hit)] By pouring their derision Upon anything we did Exposing every weakness However carefully hidden by the kids

But in town it was well known When they got home at night, their fat Psychopathic wives would thrash them Within inches of their lives

Another Brick in the Wall, Part 2 [Wa, Gi](Wa)

We don’t need no education We don’t need no thought control No dark sarcasm in the classroom Teacher leave them kids alone Hey! Teacher! Leave them kids alone! All in all it’s just another brick in the wall All in all you’re just another brick

in the wall

We don’t need no education We don’t need no thought control No dark sarcasm in the classroom Teachers leave them kids alone Hey! Teacher! Leave us kids alone! All in all you’re just another brick in the wall All in all you’re just another brick

in the wall [Teacher: ”Wrong, Do it again!” Teacher: ”If you don’t eat yer meat, you can’t have any pudding. How can you

have any pudding if you don’t eat yer meat?”

Teacher: ”You! Yes, you behind the bike sheds, stand still laddie!”]

Mother [Wa, Gi](Wa)

Mother do you think they’ll drop the bomb? Mother do you think they’ll like this song? Mother do you think they’ll try to break my balls? Oooooooa, Mother should I build a wall?

Mother should I run for president? Mother should I trust the government? Mother will they put me in the firing line? Oooooooa, Is it just a waste of time?

Hush now baby, baby, don’t you cry Mama’s gonna make all of your nightmares come true Mama’s gonna put all of her fears into you Mama’s gonna keep you right here under her wing She won’t let you fly but she might let you sing Mama’s gonna keep baby cosy and warm Ooooh Babe Ooooh Babe Oooooh Babe Of course mama’s gonna help build a wall

Mother do you think she’s good enough for me? Mother do you think she’s dangerous to me? Mother will she tear your little boy apart? Oooooooa, Mother will she break my heart?

Hush now baby, baby, don’t you cry Mama’s gonna check out all your girlfriends for you Mama won’t let anyone dirty get through Mama’s gonna wait up until you get in Mama will always find out where You’ve been Mama’s gonna keep baby healthy and clean Ooooh Babe Oooh Babe Oooh Babe You’ll always be baby to me

Mother, did it need to be so high?

Goodbye Blue Sky [Gi, Wa](Wa)

[Child: ”Look mummy, there’s an aeroplane up in the sky!”]

Oooooh Oooooh Oooooh Did did did did you see the frightened ones? Did did did did you hear the falling bombs? Did did did did you ever wonder Why we had to run for shelter When the promise of a brave new world Unfurled beneath the clear blue sky?

Oooooh Oooooh Did did did did you see the frightened ones? Did did did did you hear the falling bombs? The flames are all long gone But the pain lingers on

Goodbye Blue sky Goodbye Blue sky Goodbye Goodbye

[Announcer: ”The 11:15 from Newcastle is now approaching ... The 11:18 arrival ... ”]

Empty Spaces [Wa](Wa)

[(Backwards message)

Wa: ”Congratulations, You have just discovered the secret message. Please send your answer to ’Old Pink c/o The Funny Farm, Chalfont” Unknown voice: ”Roger, Caroline’s on the phone” Wa: ”Okay.”]

What shall we use to fill the empty Spaces where we used to talk? How shall I fill the final places? How should I complete the wall?

Young Lust [Gi](Wa)

I am just a new boy Stranger in this town Where are all the good times? Who’s gonna show this stranger around?

Ooooh I need a dirty woman Ooooh I need a dirty girl

Will some woman in this desert land Make me feel like a real man? Take this rock and roll refugee Oooh Babe set me free

Ooooh I need a dirty woman Ooooh I need a dirty girl Ooooh I need a dirty woman Ooooh I need a dirty girl [(Phone rings...Clunk of receiver being lifted) Man: ”Hello?” Operator: ”Yes, a collect call for Mrs. Floyd from Mr. Floyd. Will you accept

the charge from United States?”

(Clunk! of phone being put down)

Operator: ”Oh, he’s hung up! That’s your residence, right? I wonder why he hung up? Is there supposed to be someone else there besides your wife there to answer?”

(Phone rings again...clunk of receiver being picked up)

Man: ”Hello?”

Operator: ”This is United States calling, are we reaching...”

(Interrupted by phone being put down)

Operator: ”See he keeps hanging up, and it’s a man answering.”

(Whirr of connection being closed)]

One of My Turns [Wa](Wa)

[Roadie: ”Oh my God! What a fabulous room! Are all these your guitars?” (Film in background) ”I’m sorry sir, I didn’t mean to startle you!” Roadie: ”This place is bigger than our apartment!” Film: ”Let me know when you’re entering a room.” Film: ”Yes, sir!” Roadie: ”Erm, Can I get a drink of water?” Film: ”I was wondering about this” Roadie: ”You want some? Huh?” Film: ”Yes” Roadie: ”Oh wow, look at this tub! Do you wanna take a bath?” Film: ”I’ll have to find out from Mrs. Bancroft what time she wants to eat, it’s

for her maid... meal in the kitchen...” Roadie: ”What’cha watching?” Film: ”If you’ll just let me know as soon as you can when Mrs Bancroft wants

to eat...Mrs Bancroft...” Roadie: ”Hello?” Film: ”I don’t understand...” Roadie: ”Are you feeling okay?”]

Day after day, love turns grey Like the skin on a dying man And night after night, we pretend it’s all right But I have grown older and You have grown colder and Nothing is very much fun any more And I can feel one of my turns coming on I feel cold as a razor blade, Tight as a tourniquet, Dry as a funeral drum

Run to the bedroom, in the suitcase on the left You’ll find my favourite axe Don’t look so frightened This is just a passing phase One of my bad days Would you like to watch T.V.? Or get between the sheets? Or contemplate the silent freeway? Would you like something to eat? Would you like to learn to fly? Would’ya? Would you like to see me try? Aaaah no!

Would you like to call the cops? Do you think it’s time I stopped? Why are you runnin’ away?

Don’t Leave Me Now [Wa](Wa)

Ooooh Babe Don’t leave me now Don’t say it’s the end of the road Remember the flowers I sent I need you Babe To put through the shredder In front of my friends Ooooh Babe Don’t leave me now

How could you go? When you know how I need you (need you...) To beat to a pulp on a Saturday night Ooooh Babe Don’t leave me now

How can you treat me this way? Runnin’ away Oooooh Babe! Why are you runnin’ away?

Another Brick in the Wall, Part 3 [Wa](Wa)

[(Sound of many T.V.’s coming on, all on different channels)

Sports commentator: ”...The Bulls are already out there...” Pink: ”Aaaaaaaaaaaaaaaaaaaaaaaarrrrrgh!” (Smashes T.V. sets) Commercial(?): ”This Roman Meal bakery thought you’d like to know...”]

I don’t need no arms around me And I don’t need no drugs to calm me I have seen the writing on the wall Don’t think I need anything at all No! Don’t think I’ll need anything at all

All in all it was all just bricks in the wall All in all you were all just bricks in the wall

Goodbye Cruel World [Wa](Wa)

Goodbye cruel world I’m leavin’ you today Goodbye Goodbye Goodbye

Goodbye all you people There’s nothin’ you can say To make me change My mind Goodbye

Hey You [Gi, Wa](Wa)

Hey you! out there in the cold Getting lonely, getting old, can you feel me? Hey you! standing in the aisles With itchy feet and fading smiles, can you feel me? Hey you! don’t help them to bury the light Don’t give in without a fight

Hey you! out there on your own Sitting naked by the phone, would you touch me? Hey you! with your ear against the wall Waiting for someone to call out, would you touch me? Hey you! would you help me to carry the stone? Open your heart, I’m coming home

But it was only fantasy The wall was too high, as you can see No matter how he tried he could not break free And the worms ate into his brain

Hey you! out there on the road Always doing what you’re told, can you help me? Hey you! out there beyond the wall, Breaking bottles in the hall, can you help me? Hey you! don’t tell me there’s no hope at all Together we stand, divided we fall (we fall...)

[(Click of TV being turned on; dialogue from U.S. Western serial ’Gunsmoke’ is heard [episode: ’Fandango’])

Man: ”Well, only got an hour of daylight left. Better get started.” Margaret: ”Isn’t it unsafe to travel at night?” Man: ”It’ll be a lot less safe to stay here. You’re father’s gonna pick up our trail

before long.”

Margaret: ”Can Loca ride?” Loca: ”Yeah, I can ride...” Man: ”Margaret, time to go!” Chenga: ”Margaret, thank you for everything.” Margaret: ”Goodbye Chenga” Chenga: ”Goodbye miss...” Margaret: ”I’ll be back”]

Is There Anybody Out There? [Wa](Wa)

Is there anybody out there? Is there anybody out there? Is there anybody out there? Is there anybody out there?

Nobody Home [Wa](Wa)

[(Unknown T.V. show)

Man #1: ”Alright, I’ll take care of them part of the time, but there’s somebody else that needs taking care of in Washington.”

Man #2: ”Who’s that?” Man #1: ”Rose Pilchitt!” Man #2: ”Rose Pilchitt? Who’s that?” Pink: ”Shut Up!” Man #1: ”36-24-36.” (laughter) ”Does that answer your question?” Pink: ”Oi! I’ve got a little black book with me poems in!” Man #2: ”Who’s she?” Man #1: ”She was ’Miss Armoured Division’ in 1961...”]

I’ve got a little black book with my poems in Got a bag with a toothbrush and a comb in When I’m a good dog they sometimes throw me a bone in

I got elastic bands keepin’ my shoes on Got those swollen hand blues I got thirteen channels of shit on the T.V. to choose from I’ve got electric light And I’ve got second sight I’ve got amazing powers of observation And that is how I know When I try to get through On the telephone to you There’ll be nobody home

I’ve got the obligatory Hendrix perm And the inevitable pinhole burns All down the front of my favourite satin shirt I’ve got nicotine stains on my fingers I’ve got a silver spoon on a chain Got a grand piano to prop up my mortal remains I’ve got wild staring eyes And I’ve got a strong urge to fly But I got nowhere to fly to (fly to...) Ooooh Babe when I pick up the phone

[(From Gomer Pyle show) Man: ”Surprise, surprise, surprise”]

There’s still nobody home I’ve got a pair of Gohills boots and I got fading roots [(From The ’Battle of Britain’ movie) Man: ”Where the hell are you?” Radio: ”Over 47 German planes were destroyed with the loss of only 15 of our

own aircraft”

Man: ”Where the hell are you Tyler?” (Machine gun sound, followed by plane crashing)]

Vera [Wa](Wa)

Does anybody here remember Vera Lynn? Remember how she said that We would meet again Some sunny day?

Vera! Vera! What has become of you? Does anybody else in here Feel the way I do?

Bring the Boys Back Home [Wa](Wa)

Bring the boys back home Bring the boys back home Don’t leave the children on their own Bring the boys back home

[Teacher: ”Wrong! Do it again!” Man : ”Time to go!” (knocks on the door) Roadie: ”Are you feeling okay?” Operator: ”There’s a man answering, but he keeps hanging up!”]

Is there anybody out there?

Comfortably Numb [Wa, Gi](Gi, Wa)

Hello, Is there anybody in there? Just nod if you can hear me Is there anyone home? Come on now, I hear you’re feeling down Well I can ease your pain Get you on your feet again

Relax I’ll need some information first Just the basic facts Can you show me where it hurts?

There is no pain, you are receding A distant ship smoke on the horizon You are only coming through in waves Your lips move but I can’t hear what you’re sayin’ When I was a child I had a fever My hands felt just like two balloons Now I’ve got that feeling once again I can’t explain you would not understand This is not how I am I have become comfortably numb

I have become comfortably numb

O.K. Just a little pinprick There’ll be no more aaaaaaah! But you may feel a little sick Can you stand up? I do believe it’s working, good That’ll keep you going through the show Come on it’s time to go

There is no pain, you are receding A distant ship smoke on the horizon You are only coming through in waves Your lips move but I can’t hear what you’re sayin’ When I was a child I caught a fleeting glimpse Out of the corner of my eye

I turned to look but it was gone I cannot put my finger on it now The child is grown The dream is gone I have become comfortably numb

The Show Must Go On [Gi](Wa)

Aaaah Aaaah Aaaah Ooooh Ma Oooh Pa Must the show go on? Ooooh Pa Take me home Ooooh Ma Let me go

There must be some mistake I didn’t mean to let them Take away my soul Am I too old, is it too late?

Ooooh Ma Ooooh Pa Where has the feeling gone? Ooooh Ma Ooooh Pa Will I remember the songs? The show must go on

In the Flesh! [Wa](Wa)

So ya Thought ya Might like to Go to the show To feel the warm thrill of confusion That space cadet glow I’ve got some bad news for you sunshine Pink isn’t well he stayed back at the hotel And they sent us along as a surrogate band We’re gonna find out where you fans Really stand

Are there any queers in the theater tonight? Get ’em up against the wall! (’gainst the wall!) And there’s one in the spotlight He don’t look right to me Get him up against the wall! (’gainst the wall!) And that one looks Jewish! And that one’s a coon! Who let all this riff raff into the room? There’s one smoking a joint and Another with spots! If I had my way I’d have all of you shot!

Run Like Hell [Wa, Gi](Gi, Wa)

[Left channel: ”Pink Floyd, Pink Floyd...”

Right channel: ”Hammer, Hammer...”] Run Run Run Run Run Run Run Run Run Run Run Run Run Run Run Run You better make your face up in Your favourite disguise With your button down lips and your Roller blind eyes With your empty smile And your hungry heart Feel the bile rising from your guilty past With your nerves in tatters As the cockleshell shatters And the hammers batter Down your door You better run

Run Run Run Run Run Run Run Run Run Run Run Run Run Run Run Run You better run all day And run all night Keep your dirty feelings Deep inside And if you’re Taking your girlfriend Out tonight You better park the car Well out of sight Cause if they catch you in the back seat Tryin’ to pick her locks They’re gonna send you back to mother In a cardboard box You better run! (Run, run...)

[Man: ”Hey, open up! HaHaHaHaHaaaaaaaaaa!” (Sound of car skidding, followed by loud scream) Left channel: ”Pink Floyd, Pink Floyd...” Right channel: ”Hammer, Hammer...”]

Waiting For the Worms [Wa, Gi](Wa)

[Eins, zwei, drei, a ha!] Ooooh You cannot reach me now Ooooh No matter how you try Goodbye cruel world it’s over Walk on by

Sitting in a bunker here behind my wall Waiting for the worms to come (Worms to come...) In perfect isolation here behind my wall Waiting for the worms to come (Worms to come...)

[Man with megaphone: ”Testing one two one two We’ve orders to convene at one fifteen outside Brixton Town Hall where we will be ...”]

Waiting to cut out the deadwood Waiting to clean up the city Waiting to follow the worms Waiting to put on a black shirt Waiting to weed out the weaklings Waiting to smash in their windows And kick in their doors Waiting for the final solution To strengthen the strain Waiting to follow the worms Waiting to turn on the showers And fire the ovens Waiting for the queens and the coons and the reds and the Jews Waiting to follow the worms

Would you like to see Britannia Rule again my friend? All you have to do is follow the worms Would you like to send our coloured cousins Home again my friend? All you need to do is follow the worms

[Man with megaphone: The Worms will convene outside Brixton Bus Station. We’ll be moving along at about 12 o’clock down Stockwell Road ... Abbot’s Road ... twelve minutes to three we’ll be moving along Lambeth Road towards Vauxhall Bridge. Now when we get to the other side of Vauxhall Bridge we’re in Westminster (Borough) area. It’s quite possible we may encounter some Jew boys ... by the way we go...]

Stop [Wa](Wa)

Stop! I wanna go home Take off this uniform And leave the show But I’m waiting in this cell Because I have to know Have I been guilty all this time?

The Trial [Wa](Wa, Ezrin)

Good morning Worm Your Honour The Crown will plainly show The prisoner who now stands before you Was caught red handed showing feelings Showing feelings of an almost human nature This will not do Call the schoolmaster!

I always said he’d come to no good In the end your honour If they’d let me have my way I could Have flayed him into shape But my hands were tied The bleeding hearts and artists Let him get away with murder Let me hammer him today

Crazy toys in the attic I am crazy Truly gone fishing They must have taken my marbles away Crazy toys in the attic he is crazy

You little shit, you’re in it now I hope they throw away the key You should have talked to me more often Than you did, but no! You had to Go your own way. Have you broken any Homes up lately? ”Just five minutes, Worm, Your Honour Him and me alone”

Baaaaaaaaaabe! Come to mother baby let me hold you In my arms M’lord I never wanted him to get in any trouble Why’d he ever have to leave me? Worm, Your Honour, let me take him home

Crazy over the rainbow I am crazy Bars in the window There must have been a door there in the wall When I came in Crazy over the rainbow he is crazy

The evidence before the court is Incontravertable, there’s no need for The jury to retire In all my years of judging I have never heard before of Someone more deserving The full penalty of law The way you made them suffer Your exquisite wife and mother Fills me with the urge to defecate!

[Man: ”Go on Judge! Shit on him!” (or is it ”Sit on it!”?)] Since my friend you have revealed your Deepest fear I sentence you to be exposed before Your peers Tear down the wall!

Outside the Wall [Wa](Wa)

All alone, or in two’s, The ones who really love you Walk up and down outside the wall Some hand in hand And some gathered together in bands The bleeding hearts and the artists Make their stand

And when they’ve given you their all Some stagger and fall after all it’s not easy Banging your heart against some mad bugger’s Wall

[Man: ”Isn’t this where...”]

Checked against audio and lyrics sheet (MFSL CD edition) 950620

Songs only in The Wall movie

What Shall We Do Now?

What shall we use to fill the empty spaces Where waves of hunger roar? Shall we set out across this sea of faces In search of more and more applause?

Shall we buy a new guitar? Shall we drive a more powerful car? Shall we work straight through the night? Shall we get into fights? Leave the lights on? Drop bombs? Do tours of the east? Contract diseases? Bury bones? Break up homes? Send flowers by phone? Take to drink? Go to shrinks? Give up meat? Rarely sleep? Keep people as pets? Train dogs? Race rats? Fill the attic with cash? Bury treasure? Store up leisure But never relax at all? With our backs to the wall

When the Tigers Broke Free

It was just before dawn One miserable morning in black ’forty four. When the forward commander Was told to sit tight When he asked that his men be withdrawn. And the Generals gave thanks As the other ranks held back The enemy tanks for a while. And the Anzio bridgehead Was held for the price Of a few hundred ordinary lives.

And kind old King George Sent Mother a note When he heard that father was gone. It was, I recall, In the form of a scroll, With gold leaf and all. And I found it one day In a drawer of old photographs, hidden away. And my eyes still grow damp to remember His Majesty’d signed With his own rubber stamp.

It was dark all around. There was frost in the ground When the tigers broke free. And no one survived From the Royal Fusiliers Company C. They were all left behind, Most of them dead, The rest of them dying. And that’s how the High Command Took my daddy from me.

A Collection of Great Dance Songs (23/11 1981)[37/31]

One of These Days [Ma](Wa, Wr, Gi, Ma) Money [Gi](Wa) Sheep [Wa](Wa){Raving and Drooling} Shine on You Crazy Diamond, Parts I–V; VII [Wa] (Wr, Wa, Gi; Gi, Wa, Wr; Wa, Gi, Wr; Gi, Wr, Wa; Wa; Wa, Gi, Wr) Wish You Were Here [Gi](Wa, Gi) Another Brick in the Wall, Part 2 [Wa, Gi](Wa)

the final cut (21/3 1983)[1/6]

the post war dream [Wa](Wa)

[Car sound, switching on of car radio

”...announced plans to build a nuclear fallout shelter at Peterborough in Cambridgeshire...”

phzzt! of returning ”...three high court judges have cleared the way...” phzzt! ”...It was announced today, that the replacement for the Atlantic Conveyor the

container ship lost in the Falklands conflict would be built in Japan, a spokesman for...”

phzzt! ”...moving in. They say the third world countries, like Bolivia, which produce

the drug are suffering from rising violence ... ]

tell me true tell me why was jesus crucified was it for this that daddy died? was it you? was it me? did i watch too much t.v.? is that a hint of accusation in your eyes? if it wasn’t for the nips being so good at building ships the yards would still be open on the clyde and it can’t be much fun for them beneath the rising sun with all their kids committing suicide what have we done maggie what have we done what have we done to england should we shout should we scream ”what happened to the post war dream?” oh maggie maggie what have we done?

your possible pasts [Wa](Wa)

they flutter behind you your possible pasts some brighteyed and crazy some frightened and lost a warning to anyone still in command

[Cattle truck noises ”Ranks! Fire!”] of their possible future to take care in derilict sidings the poppies entwine with

cattle trucks lying in wait for the next time do you remember me? how we used to be? do you think we should be closer? she stood in the doorway the ghost of a smile haunting her face like a cheap

hotel sign her cold eyes imploring the men in their macs for the gold in their bags or the knives in their backs stepping up boldly one put out his hand he said, ”i was just a child then now i’m only a man”

do you remember me? how we used to be? do you think we should be closer? by the cold and religious we were taken in hand shown how to feel good and

told to feel bad [tongue tied and terrified we learned how to pray now our feelings run deep and cold as the clay] strung out behind us the banners and flags of our possible pasts lie in tatters and rags

do you remember me? how we used to be? do you think we should be closer?

one of the few [Wa](Wa)

when you’re one of the few to land on your feet what do you do to make ends meet? teach make ’em mad, make ’em sad, make ’em add two and two make ’em me, make ’em you, make ’em do what you want them to make ’em laugh, make ’em cry, make ’em lie down and die

the hero’s return [Wa](Wa)

jesus jesus what’s it all about trying to clout these little ingrates into shape when i was their age all the lights went out there was no time to whine and mope about

and even now part of me flies over dresden at angels one five though they’ll never fathom it behind my sarcasm desperate memories lie

sweetheart sweetheart are you fast asleep, good that’s the only time that i can really speak to you and there is something that i’ve locked away a memory that is too painful to withstand the light of day

when we came back from the war the banners and flags hung on everyones door we danced and we sang in the street and

the church bells rang but burning in my heart the memory smoulders on of the gunners dying words on the intercom

the gunners dream [Wa](Wa)

floating down through the clouds memories come rushing up to meet me now but in the space between the heavens and the corner of some foreign field i had a dream i had a dream goodbye max goodbye ma after the service when you’re walking slowly to the car and the silver in her hair shines in the cold november air you hear the tolling bell and touch the silk in your lapel and as the tear drops rise to meet the comfort of the band you take her frail hand and hold on to the dream

a place to stay

[”Oi! A real one ...”]

enough to eat somewhere old heroes shuffle safely down the street where you can speak out loud about your doubts and fears and what’s more no-one ever disappears you never hear their standard issue kicking in your door you can relax on both sides of the tracks and maniacs don’t blow holes in bandsmen by remote control and everyone has recourse to the law and no-one kills the children anymore no-one kills the children anymore

night after night going round and round my brain his dream is driving me insane in the corner of some foreign field the gunner sleeps tonight what’s done is done we cannot just write off his final scene take heed of the dream take heed

paranoid eyes [Wa](Wa)

button your lip and don’t let the shield slip take a fresh grip on your bullet proof mask and if they try to break down your disguise with their questions you can hide hide hide

[Man: ”I’ll tell you what, I’ll give you three blacks, and play you for five ...” Man: ”Ta! You was unlucky there son” Man: ”Time gentleman!”] behind paranoid eyes you put on your brave face and slip over the road for a jar fixing your grin as

you casually lean on the bar laughing too loud at the rest of the world with the boys in the crowd you can hide hide hide behind petrified eyes

you believed in their stories of fame fortune and glory now you’re lost in a haze of alchohol soft middle age the pie in the sky turned out to be miles too high and you hide hide hide behind brown and mild eyes

[”Oi!”]

get your filthy hands off my desert [Wa](Wa)

[”Oi...Get your filthy hands off my desert!”

”What ’e say?”]

brezhnev took afghanistan and begin took beirut galtieri took the union jack and maggie over lunch one day took a cruiser with all hands apparently to make him give it back

the fletcher memorial home [Wa](Wa)

take all your overgrown infants away somewhere

and build them a home a little place of their own the fletcher memorial home for incurable tyrants and kings

and they can appear to themselves every day on closed circuit t.v. to make sure they’re still real it’s the only connection they feel ”ladies and gentlemen, please welcome reagan and haig mr. begin and friend mrs. thatcher and paisly

[”Hello Maggie!”]

mr. brezhnev and party [”Who’s the bald chap?”] the ghost of mccarthy and the memories of nixon [”Good-bye!”] and now adding colour a group of anonymous latin- american meat packing glit-

terati”

did they expect us to treat them with any respect

they can polish their medals and sharpen their smiles, and amuse themselves playing games for a while boom boom, bang bang, lie down you’re dead

safe in the permanent gaze of a cold glass eye their favourite toy they’ll be good girls and boys in the fletcher memorial home for colonial wasters of life and limb

is everyone in? are you having a nice time? now the final solution can be applied

southampton dock [Wa](Wa)

they disembarked in 45 and no one spoke and no one smiled there were too many spaces in the line and gathered at the cenotaph all agreed with hand on heart to sheath the sacrificial knifes

but now

she stands upon southampton dock with her handkerchief and her summer frock clings to her wet body in the rain in quiet desperation knuckles white upon the slippery reins she bravely waves the boys goodbye again

still the dark stain spreads between their shoulder blades a mute reminder of the poppy fields and graves and when the fight was over we spent what they had made but in the bottom of our hearts we felt the final cut

the final cut [Wa](Wa)

through the fish eyed lens of tear stained eyes i can barely define the shape of this moment in time and far from flying high in clear blue skies i’m spiralling down to the hole in the ground where i hide

if you negotiate the minefield in the drive and beat the dogs and cheat the cold electronic eyes and if you make it past the shotguns in the hall dial the combination open the priesthole and if i’m in i’ll tell you [what’s behind the wall]

there’s a kid who had a big hallucination making love to girls in magazines he wonders if you’re sleeping with your new found faith could anybody love him or is it just a crazy dream

and if i show you my dark side will you still hold me tonight and if i open my heart to you show you my weak side what would you do would you sell your story to rolling stone would you take the children away and leave me alone and smile in reassurance as you whisper down the phone would you send me packing or would you take me home

thought i oughta bare my naked feelings thought i oughta tear the curtain down i held the blade in trembling hands prepared to make it but just then the phone rang i never had the nerve to make the final cut

[”Hello? Listen, I think I’ve got it. Okay, listen its a HaHa!”]

not now john [Gi,Wa](Wa)

fuck all that we’ve got to get on with these got to compete with the wily japanese there’s too many home fires burning and not enough trees so fuck all that we’ve got to get on with these

cant stop lose job mind gone silicon ? what bomb get away pay day make hay break down need fix big six clickity click hold on oh no brrrrrrrrrring bingo!

make em laugh make em cry make em dance in the aisles make em pay make em stay make em feel ok

not now john we’ve got to get on with the film show hollywood waits at the end of the rainbow who cares what it’s about as long as the kids go so not now john we’ve got to get on with the show

hang on john i’ve got to get on with this don’t know what it is but it fits on here like ........... come at the end of the shift we’ll go and get pissed but not now john i’ve got to get on with this

hold on john i think there’s something good on i used to read books but ............ could be the news or some other abuse or it could be reusable shows

fuck all that we’ve got to get on with these got to compete with the wily japanese no need to worry about the vietnamese got to bring the russian bear to his knees well, maybe not the russian bear maybe the swedes we showed argentina now let’s go and show these make us feel tough and won’t maggie be pleased nah nah nah nah nah nah! wow! one two three four!

s’cusi dove il bar

[”What?”]

se para collo pou eine toe bar s’il vous plait ou est le bar

[” ... say it in english!”]

oi’ where’s the fucking bar john! [” ... oh, now you’re talking!”] go! go maggie! [” ... Hammer, Hammer, Hammer, Hammer, now!”]

two suns in the sunset [Wa](Wa)

in my rear view mirror the sun is going down sinking behind bridges in the road i think of all the good things that we have left undone and i suffer premonitions

confirm suspicions of the holocaust to come

the rusty wire that holds the cork that keeps the anger in gives way and suddenly it’s day again the sun is in the east even though the day is done two suns in the sunset could be the human race is run

like the moment when the brakes lock and you slide towards the big truck [”Oh no!” ”Daddy, Daddy!”] you stretch the frozen moments with your fear and you’ll never hear their voices

and you’ll never see their faces you have no recourse to the law anymore

and as the windshield melts and my tears evaporate leaving only charcoal to defend finally i understand the feelings of the few ashes and diamonds foe and friend we were all equal in the end

[”...and now the weather. Tomorrow will be cloudy with scattered showers spreading from the east ... with an expected high of 4000 degrees Celsius”]

Checked against audio and lyrics sheet (Japanese Sony CD edition) 950730

A Momentary Lapse of Reason (8/9 1987)[3/3]{Of Promises Broken, Signs of Life, Delusions of Maturity}

Signs of Life [instr, Gi](Gi, Ezrin)

[When the child like view of the world went, nothing replaced it ... nothing replaced it ... nothing replaced it ... I do not like being asked to ... I do not like being asked to ... I do not like being asked to ...

Other people replaced it Someone who knows]

Learning to Fly [Gi](Gi, Moore, Ezrin, Carin)

Into the distance, a ribbon of black Stretched to the point of no turning back A flight of fancy on a windswept field Standing alone my senses reeled A fatal attraction holding me fast, how Can I escape this irresistible grasp?

Can’t keep my eyes from the circling skies Tongue-tied and twisted Just an earth-bound misfit, I

Ice is forming on the tips of my wings Unheeded warnings, I thought I thought of everything No navigator to guide my way home Unladened, empty and turned to stone

A soul in tension that’s learning to fly Condition grounded but determined to try Can’t keep my eyes from the circling skies Tongue-tied and twisted just an earth-bound misfit, I

[Conversation between Ma and the tower, during one of his first solo flights: Friction lock - set. Mixture - rich Propellers - fully forward Flaps - set - 10

degrees Engine gauges and suction - check

Mixture set to maximum percent - recheck Flight instruments... Altimeters - check both (garbled word) - on Navigation lights - on Strobes - on (to tower): Confirm 3-8-Echo ready for departure (tower): Hello again, this is now 129.4 (to tower): 129.4. It’s to go. (tower): You may commence your takeoff, winds over 10 knots. (to tower): 3-8-Echo Easy on the brakes. Take it easy. Its gonna roll this time. Just hand the power gradually, and it... ]

Above the planet on a wing and a prayer, My grubby halo, a vapour trail in the empty air, Across the clouds I see my shadow fly Out of the corner of my watering eye A dream unthreatened by the morning light Could blow this soul right through the roof of the night

There’s no sensation to compare with this Suspended animation, A state of bliss Can’t keep my eyes from the circling skies Tongue-tied and twisted just an earthbound misfit, I

The Dogs of War [Gi](Gi, Moore)

Dogs of war and men of hate With no cause, we don’t discriminate Discovery is to be disowned Our currency is flesh and bone Hell opened up and put on sale Gather ’round and haggle For hard cash, we will lie and deceive Even our masters don’t know the web we weave

One world, it’s a battleground One world, and we will smash it down One world ... One world

Invisible transfers, long distance calls, Hollow laughter in marble halls Steps have been taken, a silent uproar Has unleashed the dogs of war You can’t stop what has begun Signed, sealed, they deliver oblivion We all have a dark side, to say the least And dealing in death is the nature of the beast

One world, it’s a battleground One world, and we will smash it down One world ... One world

The dogs of war don’t negotiate The dogs of war won’t capitulate, They will take and you will give, And you must die so that they may live You can knock at any door, But wherever you go, you know they’ve been there before Well winners can lose and things can get strained But whatever you change, you know the dogs remain.

One world, it’s a battleground One world, and we will smash it down One world ... One world

One Slip [Gi](Gi, Manzanera)

A restless eye across a weary room A glazed look and I was on the road to ruin The music played and played as we whirled without end No hint, no word her honour to defend

I will, I will she sighed to my request And then she tossed her mane while my resolve was put to the test Then drowned in desire, our souls on fire I lead the way to the funeral pyre And without a thought of the consequence I gave in to my decadence

One slip, and down the hole we fall It seems to take no time at all A momentary lapse of reason That binds a life for life A small regret, you won’t forget, There’ll be no sleep in here tonight

Was it love, or was it the idea of being in love? Or was it the hand of fate, that seemed to fit just like a glove? The moment slipped by and soon the seeds were sown The year grew late and neither one wanted to remain alone

One slip, and down the hole we fall It seems to take no time at all A momentary lapse of reason That binds a life for life A small regret, you won’t forget, There’ll be no sleep in here tonight One slip ... one slip

On the Turning Away [Gi](Gi, Moore)

On the turning away From the pale and downtrodden And the words they say Which we won’t understand ”Don’t accept that what’s happening Is just a case of others’ suffering Or you’ll find that you’re joining in The turning away”

It’s a sin that somehow Light is changing to shadow And casting it’s shroud Over all we have known Unaware how the ranks have grown Driven on by a heart of stone We could find that we’re all alone In the dream of the proud

On the wings of the night As the daytime is stirring Where the speechless unite In a silent accord Using words you will find are strange And mesmerised as they light the flame Feel the new wind of change On the wings of the night

No more turning away From the weak and the weary No more turning away From the coldness inside Just a world that we all must share It’s not enough just to stand and stare Is it only a dream that there’ll be No more turning away?

Yet Another Movie [Gi](Gi, Leonard)

One sound, one single sound One kiss, one single kiss A face outside the window pane However did it come to this?

A man who ran, a child who cried A girl who heard, a voice that lied The sun that burned a fiery red The vision of an empty bed

The use of force, he was so tough She’ll soon submit, she’s had enough The march of fate, the broken will Someone is lying very still

He has laughed and he has cried He has fought and he has died He’s just the same as all the rest He’s not the worst, he’s not the best

And still this ceaseless murmuring The babbling that I brook The seas of faces, eyes upraised The empty screen, the vacant look

A man in black on a snow white horse, A pointless life has run its course, The red rimmed eyes, the tears still run As he fades into the setting sun

[Dialogue from ’Casablanca’:

Rick: You’re getting on that plane with Victor where you belong.

Ilsa: But Richard, no, I, I - -

Rick: You’ve got to listen to me. Do you have any idea what you’ve have to look forward to if you stayed here? Nine chances out of ten we’d both wind up in a concentration camp. Isn’t that true, Louis?

Renault: I’m afraid Major Strasser would insist.

Ilsa: You’re saying this only to make me go.

Rick: I’m saying it because it’s true. Inside of us we both know you belong with Victor. You’re part of his work, the thing that keeps him going ... if that plane leaves the ground and you’re not with him, you’ll regret it.

Ilsa: No.

Rick: Maybe not today, maybe not tomorrow, but soon, and for the rest of your life.

Ilsa: But what about us? ... But what about us? ... But what about us?]

Round and Round [instr](Gi)

A New Machine, Part I [Gi](Gi)

I have always been here I have always looked out from behind the eyes It feels like more than a lifetime Feels like more than a lifetime

Sometimes I get tired of the waiting Sometimes I get tired of being in here Is this the way it has always been? Could it ever have been different?

Do you ever get tired of the waiting? Do you ever get tired of being in there? Don’t worry, nobody lives forever, Nobody lives forever

Terminal Frost [instr](Gi)

A New Machine, Part II [Gi](Gi)

I will always be in here I will always look out from behind these eyes It’s only a lifetime It’s only a lifetime It’s only a lifetime

Sorrow [Wr, Gi](Gi)

The sweet smell of a great sorrow lies over the land Plumes of smoke rise and merge into the leaden sky: A man lies and dreams of green fields and rivers, But awakes to a morning with no reason for waking

He’s haunted by the memory of a lost paradise In his youth or a dream, he can’t be precise He’s chained forever to a world that’s departed It’s not enough, it’s not enough

His blood has frozen and curdled with fright His knees have trembled and given way in the night His hand has weakened at the moment of truth His step has faltered

One world, one soul Time pass, the river rolls

And he talks to the river of lost love and dedication And silent replies that swirl invitation Flow dark and troubled to an oily sea A grim intimation of what is to be There’s an unceasing wind that blows through this night And there’s dust in my eyes, that blinds my sight And silence that speaks so much louder that words, Of promises broken

Delicate Sound of Thunder (22/11 1988)[11/11]

Shine on You Crazy Diamond, Parts I–V [Gi] (Wr, Wa, Gi; Gi, Wa, Wr; Wa, Gi, Wr; Gi, Wr, Wa; Wa) Signs of Life [instr,Gi](Gi, Ezrin) Learning to Fly [Gi](Gi, Moore, Ezrin, Carin) Yet Another Movie [Gi](Gi, Leonard) Round and Round [instr](Gi) Sorrow [Wr, Gi](Gi) The Dogs of War [Gi](Gi, Moore) On the Turning Away [Gi](Gi, Moore) One of These Days [Ma](Wa, Wr, Gi, Ma) Time [Gi, Wr](Ma, Wa, Wr, Gi) On the Run [instr](Gi, Wa) The Great Gig in the Sky [M. Taylor, R. Fury](Wr){The Mortality Sequence; Ecclesiastics} Wish You Were Here [Gi](Wa, Gi) Us and Them [Gi, Wr](Wa, Wr){The Violent Sequence} Money [Gi](Wa) Another Brick in the Wall, Part 2 [Gi](Wa) Comfortably Numb [Gi, Jon Carin, Wr](Gi, Wa) One Slip [Gi](Gi, Manzanera) Run Like Hell [Guy Pratt, Gi](Gi, Wa)

The  Division  Bell

(30/3 1994)[1/1]{Awaken to the Sense of Reality}

Cluster One [instr](Wr, Gi)

[Crackling noises beneath the earth’s crust, recorded by California seismology buff G. William Forgey]

What Do You Want From Me [Gi, Wr](Gi, Wr, Samson)

As you look around this room tonight Settle in your seat and dim the lights Do you want my blood, do you want my tears What do you want What do you want from me Should I sing until I can’t sing any more Play these strings until my fingers are raw You’re so hard to please What do you want from me

Do you think I know something you don’t know What do you want from me If I don’t promise you the answers would you go What do you want from me Should I stand out in the rain Do you want me to make a daisy chain for you I’m not the one you need What do you want from me

You can have anything you want You can drift, you can dream, even walk on water Anything you want

You can own everything you see Sell your soul for complete control Is that really what you need

You can lose yourself this night See inside there is nothing to hide Turn and face the light

What do you want from me

Poles Apart [Gi, Wr](Gi, Samson, Laird-Clowes)

Did you know … it was all going to go so wrong for you And did you see it was all going to be so right for me Why did we tell you then You were always the golden boy then And that you’d never lose that light in your eyes

Hey you … did you ever realise what you’d become And did you see that it wasn’t only me you were running from Did you know all the time but it never bothered you anyway Leading the blind while I stared out the steel in your eyes

The rain fell slow, down on all the roofs of uncertainty I thought of you and the years and all the sadness fell away from me And did you know …

I never thought that you’d lose that light in your eyes

Marooned [instr](Wr, Gi){The Whale Song} A Great Day For Freedom [Gi, Wr](Gi, Samson)

On the day the wall came down They threw the locks onto the ground And with glasses high we raised a cry for freedom had arrived

On the day the wall cane down The Ship of Fools had finally run aground Promises lit up the night like paper doves in flight

I dreamed you had left my side No warmth, not even pride remained And even though you needed me It was clear that I could not do a thing for you

Now life devalues day by day As friends and neighbours turn away And there’s a change that, even with regret, cannot be undone

Now frontiers shift like desert sands While nations wash their bloodied hands Of loyalty, of history, in shades of gray

I woke to the sound of drums The music played, the morning sun streamed in I turned and I looked at you And all but the bitted residue slipped away … slipped away

Wearing The Inside Out [Wr, Gi](Wr, Moore)

From morning to night I stayed out of sight Didn’t recognise what I’d become No more than alive I’d barely survive In a word … overrun

Won’t hear a sound From my mouth I’ve spent too long On the inside out My skin is cold To the human touch This bleeding heart’s Not beating much

I murmured a vow of silence and now I don’t even hear when I think aloud Extinguished by light I turn on the night Wear its darkness with an empty smile

I’m creeping back to life My nervous system all away I’m wearing the inside out Look at him now He’s paler somehow But he’s coming round He’s starting to

choke It’s been so long since he spoke Well he can have the words right from my mouth

And with these words I can see Clear through the clouds that covered me Just give it time then speak my name Now we can hear ourselves again

I’m holding out For the day When all the clouds Have blown away I’m with you now Can speak your name Now we can hear Ourselves again

He’s curled into the corner But still the screen is flickering With an endless stream of garbage to … curse the place In a sea of random images The selfdestructing animal Waiting for the waves to break

He’s standing on the threshold Caught in fiery anger And hurled into the furnace he’ll … curse the place He’s torn in all directions And the screen is still flickering Waiting for the flames to break

Take It Back [Gi](Gi, Ezrin, Samson, Laird-Clowes)

Her love rains down on me easy as the breeze I listen to her breathing it sounds like the waves on the sea I was thinking all about her, burning with rage and desire We were spinning into darkness; the earth was on fire

She could take it back, might take it back some day

So I spy on her, I lie to her, I make promises I cannot keep Then I hear her laughter rising, rising from the deep And I make her prove her love to me, I take all that I can take And I push her to the limit to see if she will break

She might take it back, she could take it back some day

Now I have seen the warnings, screaming from all sides It’s easy to ignore them and G-d knows I’ve tried All of this temptation, it turned my faith to lies Until I couldn’t see the danger or hear the rising tide

[Ring-a-ring o’ roses A pocket full of posies A-tishoo! A-tishoo We all fall down.] She can take it back, she will take it back some day She can take it back, she will take it back some day She will take it back, she will take it back some day

Coming Back To Life [Gi](Gi)

Where were you when I was burned and broken While the days slipped by from my window watching Where were you when I was hurt and I was helpless Because the things you say and the things you do surround me While you were hanging yourself on someone else’s words Dying to believe in what you heard I was staring straight into the shining sun

Lost in thought and lost in time While the seeds of life and the seeds of change were planted Outside the rain fell dark and slow While I pondered on this dangerous but irresistible pasttime I took a heavenly ride through our silence I knew the moment had arrived For killing the past and coming back to life

I took a heavenly ride through our silence I knew the waiting had begun And headed straight … into the shining sun

Keep Talking [Gi, Hawking](Gi, Wr, Samson)

[(Spoken:) For millions of years mankind lived just like animals Then something happened which unleashed the powers of our imagination We learned to talk.]

There’s a silence surrounding me I can’t seem to think straight I’ll sit in the corner No one can bother me I think I should speak now Why won’t you talk to me I can’t seem to speak now You never talk to me My words won’t come out right What are you thinking I feel like I’m drowning What are you feeling I’m feeling weak now You never talk to me But I can’t show my weakness What are you thinking I sometimes wonder What are you feeling Where do we go from here

[(Spoken:) It doesn’t have to be like this All we need to do is make sure we keep talking.]

Why won’t you talk to me I feel like I’m drowning You never talk to me You know I can’t breathe now What are you thinking We’re going nowhere What are you feeling We’re going nowhere Why won’t you talk to me You never talk to me What are you thinking Where do we go from here

[(Spoken:) It doesn’t have to be like this All we need to do is make sure we keep talking.]

Lost For Words [Gi](Gi, Samson)

I was spending my time in the doldrums I was caught in a cauldron of hate I felt persecuted and paralysed I thought that everything else would just wait

While you are wasting your time on your enemies Engulfed in a fever of spite Beyond your tunnel vision reality fades Like shadows into the night

To martyr yourself to caution Is not going to help at all Because there’ll be no safety in numbers When the Right One walks out of the door

Can you see your days blighted by darkness? Is it true you beat your fists on the floor? Stuck in a word of isolation While the ivy grows over the door

So I open my door to my enemies And I ask could we wipe the slate clean But they tell me to please go fuck myself You know you just can’t win

High Hopes [Gi](Gi, Samson)

Beyond the horizon of the place we lived when we were young In a world of magnets and miracles Our thoughts strayed constantly and without boundary The ringing of the division bell had begun

Along the Long Road and on down the Causeway Do they still meet there by the Cut

There was a ragged band that followed in our footsteps Running before times took our dreams away Leaving the myriad small creatures trying to tie us to the ground To a life consumed by slow decay

The grass was greener The light was brighter When friends surrounded The nights of wonder

Looking beyond the embers of bridges glowing behind us To a glimpse of how green it was on the other side Steps taken forwards but sleepwalking back again Dragged by the force of some sleeping tide

At a higher altitude with flag unfurled We reached the dizzy heights of that dreamed of world

Encumbered forever by desire and ambition There’s a hunger still unsatisfied Our weary eyes still stray to the horizon Though down this road we’ve been so many times

The grass was greener The light was brighter The taste was sweeter The nights of wonder With friends surrounded The dawn mist glowing The water flowing The endless river

Forever and ever

[(At 8:18:) Phone conversation:

Stephen O’Rourke: ”Hello?”

Charlie: ”Yeah...”

SO’R: ”Is that Charlie?”

C: ”Yes.”

SO’R: ”Hello Charlie.”

(C mumbles something and hangs up)

S’OR: ”Great!” (hangs up)]

Singles

Arnold Layne [Ba](Ba)

Arnold Layne had a strange hobby Collecting clothes Moonshine washing line They suit him fine On the wall hung a tall mirror Distorted view, see through baby blue Oh, Arnold Layne It’s not the same, takes two to know Two to know, two to know Why can’t you see? Arnold Layne, Arnold Layne, Arnold Layne, Arnold Layne

Now he’s caught - a nasty sort of person. They gave him time Doors bang - chain gang - he hates it Oh, Arnold Layne It’s not the same, takes two to know Two to know, two to know Why can’t you see? Arnold Layne, Arnold Layne, Arnold Layne, Arnold Layne Don’t do it again!

Candy and a Currant Bun [Ba](Ba){Let’s Roll Another One}

Oh my, girl sitting in the sun Go buy, candy and a currant bun I like, to see you run Like that......

Oooh, don’t talk to me Please just walk with me Please you know I’m feeling frail

It’s true, sun shining very bright It’s you, who I’m gonna love tonight Ice cream, taste good in the afternoon Ice cream, taste good if you eat it soon

Oooh, don’t touch me child Please you know you drive me wild Please you know I’m feeling frail

Don’t try another cat Don’t go where other you must know why Very very very frail

Oh my, girl sitting in the sun Go buy, candy and a currant bun I like, to see you run Like that......

See Emily Play [Ba](Ba){Games for May}

Emily tries but misunderstands, ah ooh She often inclined to borrow somebody’s dreams till tomorrow There is no other day Let’s try it another way You’ll lose your mind and play Free games for May See Emily play Soon after dark Emily cries, ah ooh Gazing through trees in sorrow hardly a sound till tomorrow There is no other day Let’s try it another way You’ll lose your mind and play Free games for May See Emily play Put on a gown that touches the ground, ah ooh Float on a river forever and ever, Emily There is no other day Let’s try it another way You’ll lose your mind and play Free games for May See Emily play

Apples and Oranges [Ba](Ba)

Got a flip-top pack of cigarettes in her pocket Feeling good at the top Shopping in sharp shoes Walking in the sunshine town feeling very cool But the butchers and the bakers in the supermarket stores Getting everything she wants from the supermarket stores Apples and oranges Apples and oranges

Cornering neatly she trips up sweetly To meet the people She’s on time again And then I catch her by the eye then I stop and have to think What a funny thing to do ’cause I’m feeling very pink Apples and oranges Apples and oranges

I love she She loves me See you See you

Thought you might to know I’m the lorry driver man She’s on the run Down by the river side feeding ducks by the afternoon tide (quack quack) Apples and oranges Apples and oranges Apples and oranges

Paint Box [Wr](Wr)

Last night I had too much to drink Sitting in a club with so many fools Playing to rules Trying to impress but feeling rather empty I had another drink Drink - a - drink - a - drink - a - drink

What a way to spend that evening They all turn up with their friends Playing the game But in the scene I should have been Far away Away - away - away - away - away Getting up, I feel as if I’m remembering this scene before I open the door to an empty room Then I forget

The telephone rings and someone speaks She would very much like to go out to a show So what can I do - I can’t think what to say She sees through anyway Away - away - away - away - away

Out of the front door I go Traffic’s moving rather slow Arriving late, there she waits Looking very angry, as cross as she can be Be - a - be - a - be - a - be - a - be Getting up, I feel as if I’m remembering this scene before I open the door to an empty room Then I forget

It Would Be So Nice [Wr](Wr)

It would be so nice It would be so nice It would be so nice To meet, sometime

Everybody wakes, and in the morning Hot tea and can’t stop yawning Pass the butter please. Have you ever read the Daily Standard? Reading all about the plane that’s landed Upside down?

And no one knows what I did today There can be no other way But I would just like to say

It would be so nice It would be so nice It would be so nice To meet sometime It would be so nice It would be so nice It would be so nice To meet sometime

Everybody cares about the weather And everybody should know better What a waste of time Everybody lives beneath the ceiling Living out a dream that sends them reeling To a distant place

But no one knows what I did today There can be no other way But I would just like to say

It would be so nice It would be so nice It would be so nice To meet sometime It would be so nice It would be so nice It would be so nice To meet sometime It would be so nice It would be so nice It would be so nice To meet sometime

Julia Dream [Wa](Wa){Doreen’s Dream}

Sunlight bright upon my pillow Lighter than an eiderdown Will she let the weeping willow Wind his branches round Julia dream, dreamboat queen, queen of all my dreams Every night I turn the light out

Waiting for the velvet bride Will the scaly armadillo Find me where I’m hiding Julia dream, dreamboat queen, queen of all my dreams Will the misty master break me Will the key unlock my mind Will the following footsteps catch me Am I really dying Julia dream, dreamboat queen, queen of all my dreams

Point Me at the Sky [Wa, Gi](Wa)

Hey, Eugene, This is Henry McClean And I’ve finished my beautiful flying machine And I’m ringing to say That I’m leaving and maybe You’d like to fly with me And hide with me, baby

Isn’t it strange How little we change Isn’t it sad we’re insane Playing the games that we know and in tears The games we’ve been playing for thousands and thousands and ....

Pointing to the cosmic glider Pull this plastic glider higher Light the fuse and stand right back

He cried “This is my last goodbye!”

Point me at the sky and tell it fly Point me at the sky and tell it fly Point me at the sky and tell it fly

And if you survive ’til two thousand and five I hope you’re exceedingly thin For if you are stout you will have to breathe out While the people around you breathe in

People pressing on might say It’s something that I hate to say I’m slipping down to eat the ground A little refuge on my brain

Point me at the sky and tell it fly Point me at the sky and tell it fly Point me at the sky and tell it fly

And all we’ve got to say to you is goodbye It’s time to go, better run and get your bags, it’s goodbye Nobody cry, it’s goodbye Crash, crash, crash, crash, goodbye......

Careful With That Axe, Eugene [instr, Wa](Wa, Wr, Ma, Gi){Keep Smiling People; Beset by Creatures of the Deep (as part 2 of ’The Journey’); Come in Number 51, Your Time Is Up}

Unreleased songs

Scream Thy Last Scream [Ba](Ba)

Scream thy last scream old woman with a casket Blam blam your pointers point your pointers What’ll be to crunch in your sisters She’ll be scrubbing bubbles on all fours

Scream thy last scream old woman with a casket Fling your arms madly old lady with a daughter Flack chap chau fauses, mouses, houses Flittin and hittin and fittin quack quack

Watching the teley till all hours big time Fling your arms madly old lady with a daughter

Scream thy last scream old woman with a casket Blam blam your pointers point your pointers What’ll be coming to crunch in your sisters She’ll be scrubbing bubbles on all fours

Scream thy last scream old woman with a casket

Embryo [Gi,Wr](Wa)^{1}

All this love Is all I am A ball is all I am I’m so new Compared to you And I am very small Warm glow Moon bloom Always need a little more room Waiting here Seems like years Never seen the light of day

All around I hear strange sounds Come gurgling in my ear Red the light And dark the night I feel my dawn is near Warm glow Moon gloom Always need a little more room Whisper low Here I go I will see the sunshine show

Hollywood [instr]

Theme (Beat Version) [instr] Seabirds (More movie) [Wa](Wa) Fingal’s Cave (Zabriskie Point outtake) [instr](Wa,Ma,Gi,Wr) OneOne (Zabriskie Point outtake) [instr](Wa,Ma,Gi,Wr) Rain in the Country (Zabriskie Point outtake) [instr](Wa,Ma,Gi,Wr) Mademoiselle Knobs [instr] Vegetable Man [Ba](Ba)

In yellow shoes I get the blues, So I walk the street with my plastic feet With blue velvet trousers make me feel pink There’s a kind of stink about blue velvet trousers In my paisley shirt, I look a jerk, And my turquoise waistcoat is quite outta

sight. But oh, oh, my haircut looks so bad.... Vegetable man! Where are you?

So I change my gear, and I bugger my knees and I cover them up with the latest cuts My pants and socks are all in a box It does take long to find darn old socks The watch, black watch, my watch with a black face and a date in a little hole

and all the luck, its what I got, Its what I wear, Its what you see, It must be me, Its what I am! Vegetable man! Where are you?

Ah, ah ah ah, ah ah ah Hah, ah ah ah, ah ah ah - oh! I’ve been looking all over the place for a place for me But it ain’t anywhere It just ain’t anywhere.

Vegetable man, Vegetable man, Vegetable man. Vegetable man, Vegetable man, Vegetable man,

He’s the kind of fella you just gotta see if you can, Vegetable man.

The Pros and Cons of Hitchhiking (1984)

4.30 AM (Apparently They Were Travelling Abroad)

[Welshman: ”Apparently they were travelling abroad and they picked up some hitchhikers”

Surgeon: ”Drill” Man: ”Oh god!” Wife: ”Wake up you’re dreaming” Man: ”What?” Wife: ”You’re dreaming” Man: ”We were moving away from the border” Wife: ”What border? Go back to sleep” Hells Angel: ”Have a nice day” Wife: ”Huh” Hells Angel: ”Have a nice day”]

We were moving away from the border Looking for somewhere to sleep The two of us sharing the driving Two hitchhikers slumped in the back seat

[Hitchhiker: ”Hello”]

1Unfinished version released on “Works” 1983

I sneaked a quick look in the mirror She gave me a smile I said: ”Is anyone hungry? Should we stop for a while?” So we pulled off into a layby Her dress blew up over her head I said: ”Would you like to come with me?” She said something foreign under her breath

And the sun shone down on her lovely young limbs I thought to myself she’s much too good for him I lay down beside her with tears in my eyes She said

4.33 AM (Running Shoes)

So I stood by the roadside The soles of my running shoes gripping the tarmac Like gunmetal magnets Fixed on the front of her Fassbinder face Was the kind of a smile That only a rather dull child could have drawn While attempting a graveyard in the moonlight But she was impressed You could see that she thought I looked fine And when she turned sweeter The reason (between you and me) was She’d just seen my green Lamborghini

[Man: ”I think he was alone”]

So we went for a spin in the country To feel the wind in our hair To feel the power of my engine To feel the thrill of desire

And then in the trees I heard a twig snap Warning lights flashed on my map I opened my eyes and to my surprise

4.37 AM (Arabs with Knives and West German Skies)

There were Arabs with knives At the foot of the bed Right at the foot of the bed

[Arab: ”Don’t cut the woman, Mohamed, I want the woman”] Oh my God how did they get in here? I thought we were safe home in England She said: ”Come on now kid, it was wrong what you did You’ve got to admit it was wrong what you did You’ve got to admit it was wrong.”

[Man: ”Oh god....jesus”] Sleep Sleep I know that I’m only dreaming

[Man: ”Leave her alone..get out..out..get out of my house”] Through closed eyes I see West German skies on the ceiling And I want to get back To the girl with the rucksack To feel her flaxen hair I want to be there See the sun going down Behind Krupps steelworks On the outskirts of some German town

[Mein host: ”Guten Abend meiner Damen und Herren Ha Ha Ha Ha Willkommen in Konigsburg Ha Ha Ha Ha Wollen zie danzen mit mir oder drinken Mehr Ha Ha Ha Ha”] Thank you but...... This young lady and I Will just finish this bottle of wine It was kind of you...but I think we’ll just say goodnight.

[Man: ”Leave us alone”] Goodnight

[Man: ”Leave us alone...” ”Could I have the keys to one four three please?”

Room clerk: ”There you are” Man: ”Thank you, goodnight” Room clerk: ”Goodnight Sir” Man: ”Hello yes I’d like to order breakfast please I’d like coffee for two, and

toast with marmalade No...marmalade”]

4.39 AM (For the First Time Today, Part 2)

For the first time today I help her naked body next to mine In this hotel overlooking the Rhine I made her mine Ooh babe...ooh babe Come with me and stay with me Please stay with me

[Wife: ”Uh...what is it?”] Stay with me Stay with me Stay with me

[Wife: ”No...I’m sleeping”] Stay with me

[Wife: ”Forget it”]

4.41 AM (Sexual Revolution)

Hey...girl Take out the dagger And let’s have a stab at the sexual revolution Hey girl Let freedom for all be our rallying call Tomorrow lets make...our new resolution Yeah, but tonight lie still While I plunder your sweet grave And remember Only the poor can be saved

Hey girl As I’ve always said I prefer your lips red Not what the good Lord made But what he intended Hey girl Don’t point the finger at me I am only a rat in a maze like you And only the dead go free So...please hold my hand As we blunder through the maze And remember Nothing can grow without rain

Don’t point Don’t point your finger at me I awoke in a fever The bedclothes were all soaked in sweat She said ”You’ve been having a nightmare And it’s not over yet” Then she picked up the doggy in the window The one with the waggly tail And she put him to bed between two bits of bread

4.47 AM (The Remains of Our Love)

I just cowered in the corner My pyjama coat over my head And she smiled as she finished her sandwich And her cold eyes fixed me to my dark history As she brushed the remains Of our love from the bed And when she had turned back the covers When all of the prayers had been read She said ”Come on over here you silly boy Before you catch your death of cold I was only joking Let’s leave behind the city grime Let’s not compete It could be fine in the country Couldn’t it though...come on let’s go” I said ”OK”

[Jade: ”Are we going to go now?”

Man: ”Where would you like to go darling?” Wife: ”Mmm...Vermont...Wyoming (yes)” Man: ”Wyoming...uh huh, Children!” Justin: ”What?” Man: ”We’re going to Wyoming Darling...Which way is Wyoming?” Wife: ”Hook a right here You’re going the wrong way” Man: ”I know that, I know children... Let’s see how many...Volvos we pass On

the way to our new life in the country ...One”

Wife: ”Jade don’t do that, that’s really negative”]

4.50 AM (Go Fishing)

As cars go by I cast my mind’s eye Over back packs on roof racks Beyond the horizon Where dream makers Working white plastic processors Invite the unwary To reach for the pie in the sky Go fishing my boy!

We set out in the spring With a trunk full of books about everything About solar devices And how nice natural childbirth is We cut down some trees And we trailed our ideals Through the forest glade We dammed up the stream And the kids cooled their heels In the fishing pool we’d made We held hands and we exchanged bands And we practically lived off the land

You adopted a fox cub Whose mother was somebody’s coat You fed him by hand And then snuggled him down In the grandfather bed while I wrote We grew our own maize And I only occasionally went into town To stock up on antibiotics And shells for the shotgun that I kept around I told the kids stories While you worked your loom And the sun went down sooner each day.

[Man: ”Chapter six in which Eeyore has a birthday And gets two presents” (Man pauses to light joint)

Jade: ”Daddy...come on dad”

Man: ”Eeyore the old grey donkey stood by the side Of the stream and he looked at himself in the water ’Pathetic’ he said, ’That’s what it is’ ’Good morning Eeyore’ said Pooh ’Oh’ said Pooh, He thought for a long time”

Then the leaves all fell down Our crops all turned brown It was over As the first snowflakes fell I realised all was not well in the camp The kids caught bronchitis The space heater ran out of diesel One weekend a friend from the East Rot his soul Stole your heard I said ”Fuck it then Take the kids back to town Maybe I’ll see you around”

[Man: ”I said: ’Go now’, she said: ’O.K.’”]

And so...leaving all our hopes and dreams To the wind and the rain Taking only our stash Left our litter and trash And set out on the road again On the road again

[Jade: ”Bye Bye Daddy, Bye Daddy”

Justin: ”You can bring Pearl she’s a darn nice girl But don’t bring Liza”]

4.56 AM (For the First Time Today, Part 1)

For the first time today I feel it’s really over You were my everyday excuse For playing deaf, dumb and blind And who’d have ever thought This was how it would end for you and me To carry my own millstone Out of the trees And I have to admit I don’t like it a bit Being left here beside this lonesome road. (Lonesome road Lonesome road)

4.58 AM (Dunroamin, Duncarin, Dunlivin)

[Truckdriver: ”Hey kid, you looking for a lift?...Get on up here How’s it going good buddy?”] I nailed ducks to the wall Kept my heart in dark ruins I built bungalows all over the hills Dunroamin, duncarin, dunlivin Took my girl to the country To sleep out under the moon Next thing she’s going crazy

[Trucker: ”Women are like that kid What the hell can you do?”] She waits for the real Mr. Right to come Gently removing her heart With his promises of real communication

[Trucker: ”I saw a program about that on TV...”] Who’s always picking up the tab Who built a bungalow for his mum and dad

[Man: ”Me...”] Who took you out to all the shows Who worked his fingers to the bone

[Man: ”Me...”] While you were asleep

[Man: ”It was me...I did”] I kept you in buttons and bows

[Man: ”Christ all those clothes”] So you could encourage this creep With his neat feet And his clean fingernails With his wise but twinkling eyes He’s a rock standing out in an ocean of doubt

[Trucker: ”Get movin’, get off the road ya Goddam faggot”] And compromise I’d like to go on with this bit of a song Describing this schmuck I’d like to go on, but I’m going to throw up

[Trucker: ”Not in my rig you don’t boy...get the hell out of here”]

5.01 AM (The Pros and Cons of Hitchhiking)

An angel on a Harley Pulls across to greet a fellow rolling stone Puts his bike up on it’s stand Leans back and then extends A scarred and greasy hand...he said [Hells Angel: ”How ya doin bro?...where ya been?...where ya goin’ ?”]

Then he takes your hand In some strange Californian handshake And breaks the bone

[Angels Moll: ”Have a nice day”] A housewife from Encino Whose husband’s on the golf course With his book of rules Breaks and makes a ’U’ and idles back To take a second look at you You flex your rod Fish takes the hook Sweet vodka and tobacco in her breath Another number in your little black book

These are the pros and cons of hitchhiking These are the pros and cons of hitchhiking

Oh babe, I must be dreaming I was standing on the leading edge The Eastern seaboard spread before my eyes ”Jump” said Yoko ”Oh No! I’m too scared and too good looking” I cried ”Go on”, she said ”Why don’t you give it a try? Why prolong the agony all men must die”

Do you remember Dick Tracy? Do you remember Shane?

[Joey: ”And mother wants you”] Could you see him selling tickets Where the buzzard circles over

[Joey: ”Shane”] The body of the plain Did you understand the music Yoko Or was it all in vain?

[Joey: ”Shane”] The bitch said something mystical ”Herro” So I stepped back on the kerb again

These are the pros and cons of hitchhiking These are the pros and cons of hitchhiking

Oh babe, I must be dreaming again

These are the pros and cons of hitchhiking

5.06 AM (Every Strangers Eyes)

[Waitress: ”Hello...you wanna cup of coffee?”

Trucker: ”Hey...turn the fucking Juke Box down”

Waitress: ”I’m sorry, would you like a cup of coffee? Ok you take cream and sugar?”] In truck stops and hamburger joints In Cadillac limousines In the company of hasbeens And bent-backs and sleeping forms On pavement steps In libraries and railway stations In books and banks In the pages of history In suicidal cavalry attacks I recognise... Myself in every stranger’s eyes

And in wheelchairs by monuments Under tube trains in commuter accidents In council care and county courts At Easter fairs in sea-side resorts In drawing rooms and city morgues In award winning photographs Of life rafts in the China seas In transit camps, under arc lamps On loading ramps In faces blurred by rubber stamps I recognise... Myself in every stranger’s eyes

And now from where I stand Upon this hill I plundered from the pool I look around, I search the skies I shade my eyes, so nearly blind And I see signs of half remembered days I hear bells that chime in strange familiar ways I recognise... The hope you kindle in your eyes

It’s oh so easy now As we lie here in the dark Nothing interferes it’s obvious How to beat the tears That threaten to snuff out The spark of our love

5.11 AM (The Moment of Clarity)

Then the moment of clarity Faded like charity does Sometimes I opened one eye And I put out my hand just to touch your soft hair To make sure in the darkness that you were still there And I have to admit I was just a little afraid, Ooooo oooh yeah But then... I had a little bit of luck You were awake I couldn’t take another moment alone.

Checked against audio and lyrics sheet (British Harvest CD edition) 950618

When the Wind Blows (/10 1986)

The Russian Missile [instr](Wa)

Towers of Faith [Wa, Clare Torrey](Wa)

Ooooh, the lonely boys In their towers of faith Ooooh, the lonely boys Locked in their towers of faith The prophet reclined In the Golan Heights Ohhh, the lonely boys He said ”This land is my land” To the Shiites Ooooh, the lonely boys And Jehova looked up from the sea of Galilee beneath He said ”I see you, you

thief! This land is my land And this sand is my sand And this band is my band”

Oh the lonely boys Lookin’ over their shoulder Checkin out every boulder in the park Where the gates are closed from hate After dark

And the Pope rolled up in his armored van He fell on his knees and kissed the land He said something that I did not understand It was in polish Then up stepped an aide, he said ”I will translate Here is what His Holiness said: ’I am the Chief Jesuit. This land is Jesus’ land. And that is all All that there is to it.’”

And in New York City The business man in his mohair suit In the world trade center Puffs on his cheroot And he said ”Well I don’t care who owns the desert sands, My brief Is with the hydrocarbons underneath”

And the sea of battle rages around the ancient tombs And mother nature licks her wounds And the lonely boys locked in their towers of faith Who are nervous in the park when the gates are closed after dark

Ooooh, the lonely boys In their towers of faith Ooooh, the lonely boys Locked in their towers of faith

Hilda’s Dream [instr, John Mills, Peggy Ashcroft](Wa)

[(Radio:) ”Radio Four, with the news at 1 o’clock this Thursday lunchtime. The Prime Minister, speaking a few minutes ago in the House of Commons, has warned that the international situation is deteriorating rapidly and that war could break out at any time in the next two or three days.”

(Man:) ”This is it! This is really it!” (Woman:) ”I shouldn’t worry too much, it’ll probably all blow over.”

The American Bomber [instr](Wa) The Anderson Shelter [instr, Mills, Ashcroft](Wa)

[(Woman:) ”It’s funny to think there’s no shelters this time. We had an old Anderson in the garden. I can see it now. We had stertians growing all over it, and we painted the front green. Painted it looked quite pretty. Next door grew cabbages on theirs.

(Man:) ”Yes, we had a Morrison. I used to sleep in it. Mmmm. I stuck pin-up girls all over the inside. Betty Grabel, Ann Shelton, Patricia Rock. The roof got all smoky, ’cause I used to read in bed with a candle.”

(Woman:) ”Yes, it was nice at the war, really. The shelters, the blackouts, cups of tea...”

(Man:) ”The A.R.P., the evacuees, London kids seeing cows for the first time....” (Woman:) ”Old Churchill on the wires...” (Man:) ”Those were the days.”]

The British Submarine [instr](Wa) The Attack [instr, Mills, Ashcroft](Wa)

[(Radio:) ”An enemy missile attack has been launched against this country. It is estimated that the missiles will arrive in approximately three minutes.” (Man:) ”God, almighty, there’s only three minutes to go!” (Woman:) ”Okay, I’ll just get the washing in.” (Man:) ”Shut up! I’m trying to listen!” (Radio:) ”Stay indoors.” (Woman:) ”Oh, dear, I’ve left the oven on.” (Man:) ”Get in! Get in! Get in!” (Woman:) ”The cake will be burned!”]

The Fallout [instr, Mills, Ashcroft](Wa)

[(Man:) ”Blimey! I...I suppose...that was it.” (Woman:) ”Wasn’t it light?” (Man:) ”Yea, terrific. You–you get terrific light with those bombs. Listen! A dog!”

(Woman:) ”What does the fallout look like, dear?” (Man:) ”Fallout? The government director neglected to mention what it would

look like. I expect it would look a bit like snow does, only gray. Very quiet. I expect they’re all are having a good lie-in after the bombing.”

(Woman:) ”Terrible smell of burning.” (Man:) ”Well, yes. Well there’s bound to be. That’s logical...” (Woman:) ”It’s like...roast meat.” (Man:) ”Yes, it does. I expect everybody will be having their Sunday dinner a bit early this week, due to the unexpected circumstances...”

(Man:) ”A cloud coming up. Looks like rain. We’ll be alright for water now for a while, my dear.” (Woman:) ”Do you think rainwater is alright to drink?” (Man:) ”Well, yes, of course it is. There’s nothing purer than rainwater, is there? Everybody knows that.”

(Woman:) ”Oh, look! My hair’s coming out.”

Hilda’s Hair [instr, Mills, Ashcroft](Wa)

(Man:) ”I shall hear no evil...and thou shalt accompany me all the days–all the days of my life...uh...lay me down in–in–in green pastures..I...I don’t know anymore...” (Woman:) ”That was nice, really. Lay me down in green pastures–” (Man:) ”Oooh yes, yes...enter the valley of the shadow of death (Woman:) ”no more, now...no more...”

Folded Flags [Wa, Carrack](Wa)

Rock a bye baby On the tree top When the wind blows The cradle will rock And babe, hate to see you fall that way Better speak to the palace than to be today

Hey Joe, where you goin’ With that gun in your hand? You can take your revenge But you’ll still feel bad There must be more to life than lucky strikes And some unlucky ones And folded flags and pipes And drums

I stood in the wings with you Our lives in the hands of a second-rate actor Holding the high ground On some old stage And babe, how do these cheated stores get so far away Will they catch you at the bottom

Hey Joe, where you goin’ With that dogma in your head? You can prove your point, But your kids will still be dead Bring down the curtain The soap opera must surely close Before the cold wind blows

Hey Joe, where you goin’ with that gun in your hand You can take your revenge But you’ll still feel bad Bring down the curtain The show must close Before the cold wind blows

So rock a bye baby On the tree top When the wind blows The cradle will rock There must be more to life than lucky strikes And some unlucky ones And folded flags and pipes And drums

Radio KAOS (1987){Home}

Radio Waves

[Rumour only sofar: ”...back to the beginning”]

[Jim: This is K.A.O.S. You and I are listening to KAOS in Los Angeles. Let’s go to the telephones now and take a request.

Billy: Hello, I’m Billy.

Jim: Yes?

Billy: I hear radio waves in my head.

Jim: You hear radio waves in your head? Ah! Is there a request that you have tonight for KAOS?]

Radio waves. Radio waves. He hears radio waves. Radio waves. The atmosphere is thin and cold The yellow sun is getting old The ozone overflows with radio waves AM, FM, weather and news Our leaders had a frank exchange of views Are you confused, radio waves.

Radio waves, radio waves AM radio waves, FM radio waves Radio waves, mindnumbing radio waves Fish-stunning radio waves Radio waves.

Magic Billy in his wheel chair Is picking up all this stuff in the air Billy is face to face with outer space Messages from distant stars The local police calling all cars, radio waves

Hear them radio waves, radio waves Jesus saves radio, radio waves Radio waves, AM radio waves, FM radio waves All them radio waves

Radio waves, radio waves, he hears radio waves Radio waves, radio waves, hopeful radio waves, dopeful radio waves Radio waves, Russian radio waves, Prussian radio waves Eastern radio waves, Western radio waves Testing radio waves, one two. One two. Radio waves. Getting through to you More code radio waves, Tobacco road radio waves South to Paloma radio waves, Oklahoma City radio waves Sitting pretty radio waves, nitty-gritty radio waves Radio waves

[Jim: Alright, that’s a song called Radio Waves. You are listening to KAOS in Los Angeles and we’ve got Billy on the line.

Billy: I’m from the valleys. Jim: You’re from the valley? Billy: No, Jim you schmuck, the Valleys; male voice choirs, Wales. Jim: Ah, you’re from Wales! Now is this sperm or blue-tip? Billy: Ha, ha, ha, ha. Very funny Jim. Jim: Sorry. Billy: Me and Benny went out. Jim: Who’s Benny?]

Who Needs Information

Me and Benny went out last night Looking for fun Supping ale in the moonlight Waiting for the dawn to come Benny pointed at a HiFi shop He said hey man look at all the stuff they’ve got How’d you make a have out of a have not Hmmmm. Who needs information When you’re working underground Just give me confirmation We could win a million pounds

Benny climbed up on a footbridge And he teetered on the parapet He said can you see the whites of their headlights Are they coming yet

Who needs information This high off the ground Just give me confirmation We could win a million pounds

Who needs information When you’re living in constant fear Just give me confirmation There’s some way out of here Some way out of here

Benny hefted a breeze block And tried to let go Got hung up on a tear drop So me and Benny went home

Who needs information When you’re living in constant fear Just give me confirmation There’s some way out of here Some way out of here

Who needs information yeah When you’re living on borrowed time Just give me confirmation There will be a winner this time

Who needs information when you’re working underground Just give me confirmation We could win a million pounds Who needs, who needs, who needs information This high off the ground Just give me confirmation We could win a million pounds - yeah

[Jim: Um. So your brother’s in jail?]

Me or Him

You wake up in the morning, get something for the pot Wonder why the sun makes the rocks feel hot Draw on the walls, eat, get laid Back in the good old days

Then some damn fool invents the wheel Listen to the whitewalls squeal You spend all day looking for a parking spot Nothing for the heart, nothing for the pot

Benny turned the dial on his Short Wave radio Oh how he wanted to talk to the people, he wanted his own show Tune in Moscow. Tune in New York Listen tot the Welsh kid talk Communicating like in the good old days

Forgive me father for I have sinned It was either me or him And a voice said Benny You fucked the whole thing up Benny your time is up Your time is up

Benny turned the dial on his Short Wave radio He wanted to talk to the people He wanted his own show Tune in Moscow. Tune in New York Listen to the Welsh kid talk communicating Like in the good old days

Forgive me Father [Welsh Policeman: Mobile One Two to Central.] For I have sinned [Welsh Policeman: We have a multiple on the A465] between Cwmbran and Cylgoch. Father it was either me or him. Father can we turn back the clock? [Welsh Policeman: Ambulance, over.] I never meant to drop the concrete block. [Welsh Policeman: Roger central, over and out.]

Benny turned the dial on his Short Wave radio He wanted to talk to the people He wanted his own show Tune in Moscow. Tune in New York Listen to the Welsh kid talk Just like in the good old days The good old days

[Radio announcer: Do you really think Iranian terrorists would have taken Americans hostage if Ronald Reagan were president? Do you really think the Russians would have invaded Afghanistan if Ronald Reagan were president? Do you really think third-rate military dictators would laugh at America and burn our flag in contempt if Ronald Reagan were president? Concerned Citizen: Well, it might work! Hostage: We as a group do most importantly want to beseech President Reagan and our fellow Americans to refrain from any form of military or violent means as an attempt, no matter how noble or heroic, to secure our freedom.

Concerned Citizen: Sure! Only it’s going to be mighty dangerous for you, Cassidy

Hoppy’s faithful sidekick: guess you don’t know Hopalong Cassidy, Mister. Adventure’s his bread, excitement’s his butter and danger, why to him that’s like strawberry jam to top it off.

Jim: This is some live rock and roll at KAOS, where rock and roll comes out of chaos and a song called ”The Powers that Be”]

The Powers That Be [Wa, Paul Carrack]

The powers that be They like a tough game No rules Some you win, some you lose Competition’s good for you They’re dying to be free They’re the powers that be They like a bomb proof cadillac Air conditioned, gold taps, Back seat gun rack, platinum hub caps They pick horses for courses They’re the market forces Nice car Jack They like order, make-up, lime light power Game shows, rodeos, star wars, TV They’re the powers that be If you see them come, You better run - run You better run on home

Sisters of mercy better join your brothers Put a stop to the soap opera right now They say the toothless get ruthless You better run on home

You better run - run You better run on home

The powers that be They like treats, tricks, carrots and sticks They like fear and loathing, they like sheep’s clothing And blacked-out vans

Blacked-out vans, contingency plans They like death or glory, they love a good story They love a good story

Sisters of mercy better join with your brothers Put a stop to the soap opera state They say the toothless get ruthless Run home before its too late You better run - run You better run on home

[Billy: Goodnight, Jim. Jim: Goodnight, Billy. Uncle David’s Great Dane: Woof, woof, woof!

The canyon - daytime. Billy plays with Great Uncle David’s Great Dane.

Paraquat Kelly: Bull heads, three red snapper, one pink snapper and your Pacific coastal trench hosemonster fish. Cynthia Fox: Ohhh! At Sky David’s juke joint of joy reports, forty under the console giggle stick ling cod, twenty-three purple perches four sledgehammerhead sharks, and what a surprise, eightyfour crabs, and no red snappers. Paraquat Kelly: Hey, and that’ll do for the triumphant return of the fish report with a beat.

Jim: We think of it as mainstreet, but to the rest of the country it’s Sunset Strip. You’re listening to KAOS in Los Angeles.]

Sunset Strip

I like staying with my Uncle Dave And I like playing with his great dane But I don’t fit I feel alien and strange Kinda outa range

I like riding in my Uncle’s car Down to the beach where the pretty girls all parade And movie stars and paparazzi play The Charles Atlas kicking sand in the face game

And I sit in the canyon with my back to the sea There’s a blood red dragon on a field of green Calling me back

Back to the Black Hills again Ooh, ooh, Billy come home Billy is searching for his native land Flicking through the stations with the dial

in his head Picking up ————– and A male voice choir on the short wave band

Billy taps out Jim’s number on the ’phone Sits shaking as he waits for Jim’s answering tone Come on my friend, speak to me please The land of my fathers is calling to me And I sit in the canyon with my back to the sea There’s a blood red dragon on a field of green Calling me back, back to the Black Hills again Ooh, ooh, Billy come home

Come on home He sits in the canyon with his back to the sea Sees a blood red dragon on a field of green He hears a male voice choir singing Billy come home Billy, Billy, come home Come on home

[Californian Weirdo: I don’t like fish. Jim: You are listening to KAOS here in Los Angeles. Californian Weirdo: I don’t like fish. Jim: Yes, we’ve established that. Ah! Do you have a request? Californian Weirdo: Shell fish, guppy, salmon, shrimp and crab and lobster, flounder.I hate fish, but I think most of all I hate fresh fish, like trout. I hate fresh trout. My least-hated, favourite fish would be sole. That way you don’t have to see the eyes. Sole has no eyes. Jim: Oh no! Californian Weirdo: I’d like to be home with my monkey and my dog Jim: Thankyou. Californian Weirdo: I’d like to be home with my monkey and my dog I’d like to be home with my monkey and my dog I’d like to be home with my monkey ... Jim: They don’t care. Shut up. Play the record.]

Home [Wa, Clare Torrey]

[Jim: Oh, God! Californian Weirdo: Sole has no eyes.]

Could be Jerusalem, or it could be Cairo Could be Berlin, or it could be Prague Could be Moscow, could be New York Could be Llanelli, and it could be Warrington Could be Warsaw, and it could be Moose Jaw Could be Rome Everybody got somewhere they call home When they overrun the defences A minor invasion put down to expenses Will you go down to the airport lounge Will you accept your second class status A nation of waitresses and waiters Will you mix their martinis Will you stand still for it Or will you take to the hills

It could be clay and it could be sand Could be desert Could be a tract of arable land Could be a house, could be a corner shop Could be a cabin by a bend in the river Could be something your old man handed down Could be something you built on your own Everybody got something he calls home

When the cowboys and Arabs draw down On each other at noon In the cool dusty air of the city boardroom Will you stand by a passive spectator Of the market dictators Will you discreetly withdraw With your ear pressed to the boardroom door Will you hear when the lion within you roars Will you take to the hills

Will you stand, will you stand for it Will you hear, ohhhh! ohhh! when the lion within you roars

Could be your father and it could be your mother Could be your sister, could be your brother Could be a foreigner, could be a Turk Could be a cyclist out looking for work. Norman Could be a king, could be the Aga khan Could be a Vietnam vet with no arms and no legs Could be a saint, could be a sinner Could be a loser or it could be a winner Could be a banker, could be a baker Could be a Laker, could be Kareem Abdul Jabar Could be a male voice choir Could be a lover, could be a fighter Could be a super heavyweight, or it could be something lighter Could be a cripple, could be a freak Could be a wop, gook, geek Could be a cop, could be a thief Could be a family of ten living in one room on relief Could be our leaders in their concrete tombs With their tinned food and their silver spoons Could be the pilot with God on his side Could be the kid in the middle of the bomb sight Could be a fanatic, could be a terrorist Could be a dentist, could be a psychiatrist Could be humble, could be proud Could be a face in the crowd Could be the soldier in the white cravat Who turns the key in spite of the fact That this is the end of the cat and mouse Who dwelt in the house Where the laughter rang and the tears were spilt The house that Jack built Where the laughter rang and the tears were spilt The house that Jack built Bang, bang, shoot, shoot White gloved thumb, Lord thy will be done He was always a good boy his mother said He’ll do his duty when he’s grown, yeah Everybody’s got someone they call home

Four Minutes [Wa, Clare Torrey]

[Billy: Four minutes and counting. Jim: O.K. Billy: They pressed the button, Jim. Jim: They pressed the button Billy, what button? Billy: The big red one. Jim: You mean THE button? Billy: Goodbye, Jim. Jim: Goodbye! Oh yes. This ain’t au revoir, it’s goodbye! Ha! Ha! Jim: This is KAOS. It’s a beautiful, balmy, Southern California summer day. It’s 80 degrees ... I said balmy ... I could say bomby ... Ha! Ha! ...O.K. I’m Jim and this is Radio KAOS and with only four minutes left to us, let’s use this as wisely as possible. Molly: Everybody got someone they call home. Jim: Out at Dodger Stadium. It’s the bottom of the seventh, the Dodgers are leading three to nothing over the Giants, and for those of you who are looking to go surfing tomorrow, too bad. (Telephone rings) Jim: I’m kinda lost in here to tell you the truth ... O.K. good. Ladies and gentlemen, if the reports that we are getting are correct, this could be it. Billy, if you’re listening to me, please call now.]

After a near miss on the plane You swear you’ll never fly again After the first kiss when you make up You swear you’ll never break up again And when you’ve just run a red light Sit shaking under the street light You swear to yourself you’ll never drink and drive again Sometimes I feel like going home You swear you’ll never let things go by again. Sometimes I miss the rain and snow And you’ll never toe the party line again And when the east wind blows Sometimes I feel like going home

[Jim: Billy, if you are listening, please call. Californian Weirdo: Sole has no eyes. Molly: Goodbye little spy in the sky. They say that cameras don’t lie. Am I happy, am I sad, am I good, am I bad? Jim: Billy, if you’re listening, please call. Californian Weirdo: Sole has no eyes, sole has no eyes Billy: Ten, nine, eight, seven Margaret Thatcher: Our own independent nuclear deterrent has helped to keep the peace. Billy: Six, five four, three, Ordinary Person: ...you’ve go a job... Billy: Two, one, Margaret Thatcher: For nearly forty years Jim: Goodbye Billy.]

The Tide is Turning (After Live Aid)

I used to think the world was flat Rarely threw my hat into the crowd I felt I had used up my quota of yearning Used to look in on the children at night In the glow of their Donald Duck light And frighten myself with the thought of my little ones burning But oh, oh, oh, the tide is turning The tide is turning

Satellite buzzing through the endless night Exclusive to moonshots and world title fights Jesus Christ imagine what it must be earning Who is the strongest, who is the best Who holds the aces, the East or the West This is the crap our children are learning But oh, oh, oh, the tide is turning The tide is turning Oh, oh, oh, the tide is turning

Now the satellite’s confused ’Cos on Saturday night The airwaves were full of compassion and light And his silicon heart warmed To the sight of a billion candles burning Oo, oo, oo, the tide is turning Oo, oo, oo, the tide is turning The tide is turning Billy

I’m not saying that the battle is won But on Saturday night all those kids in the sun Wrested technology’s sword from the hand of the War Lords Oh, oh, oh, the tide is turning The tide is turning Sylvester

The tide is turning.

”That’s it!” [(Morse Code:) Now the past is over but you are not alone Together we’ll fight Sylvester Stallone We will not be dragged down in his South China Sea of macho bullshit and mediocrity]

[Rumour only sofar: ”That’s...”]

Amused To Death (26/8 1992)

The Ballad Of Bill Hubbard [instr, Razzell]

[(switch channels) (Alf Razzell:) ”Two things that have haunted me most are the days when I had to collect the paybooks; and when I left Bill Hubbard in no man’s land.

”I was picked up and taken into their trench. And I’d no sooner taken two or three steps down the trench when I heard a call, ’Hello Razz, I’m glad to see you. This is my second night here,’ and he said ’I’m feeling bad,’ and it was Bill Hubbard, one of the men we’d trained in England, one of the original battalion. I had a look at his wound, rolled him over; I could see that it was probably a fatal wound. You could imagine what pain he was in, he was dripping with sweat; and after I’d gone about three shellholes, traversed that, had it been...had there been a path or a road I could have done better. He pummelled me, ’Put me down, put me down, I’d rather die, I’d rather die, put me down.’ I was hoping he would faint. He said ’I can’t go any further, let me die.’ I said ’If I leave you here Bill you won’t be found, let’s have another go.’ He said ’All right then.’ And the same thing happened; he couldn’t stand it any more, and I had to leave him there, in no man’s land.”

(switch channels) (girl:) ”I don’t mind about the war, that’s one of the things I like to watch, if it’s a war going on, ’cause then I know if our side’s winning, if our side’s losing...” (switch channels)]

What God Wants, Part I {Fallen Angel}

What God wants God gets God help us all What God wants God gets The kid in the corner looked at the priest And fingered his pale blue Japanese guitar The priest said God wants goodness God wants light God wants mayhem God wants a clean fight What God wants God gets Don’t look so surprised It’s only dogma The alien prophet cried The beetle and the springbok Took the Bible from its hook The monkey in the corner Wrote the lesson in his book What God wants God gets God help us all God wants peace God wants war God wants famine God wants chain stores What God wants God gets God wants sedition God wants sex God wants freedom God wants semtex What God wants God gets Don’t look so afraid I’m only joking The alien comic lied The jackass and hyena Took the feather from its hook The monkey in the corner Wrote the joke down in his book What God wants God gets God wants boarders God wants crack God wants rainfall God wants wetbacks What God wants God gets God wants voodoo God wants shrines God wants law God wants organised crime God wants crusade God wants jihad God wants good God wants bad What God wants God gets

Perfect Sense, Part I

[Backwards message: July (pause) however (pause) in the light and visions of the (very soft) issues of sterling, (pause) we have changed our minds. We have decided to include a backward message (pause) Stanley (pause) for you (pause) and all the other (dramatic pause) geek partners. (pause) (very loud, noisy, screamed sentence)]

The monkey sat on a pile of stone And he stared at the broken bone in his hand The strings of a Viennese quartet Rang out across the land And the monkey looked up at the stars And he thought to himself Memory is a stranger History is for fools And he cleaned his hands In a pool of holy writing Turned his back on the garden And set out for the nearest town Hold on hold on soldier When you add it all up The tears and the marrowbone There’s an ounce of gold And an ounce of pride in each ledger And the Germans killed the Jews And the Jews killed the Arabs And the Arabs killed the hostages And that is the news And is it any wonder That the monkey’s confused He said Mama Mama The President’s a fool Why do I have to keep reading these technical manuals And the joint chiefs of staff And the brokers on Wall Street said Don’t make us laugh You’re a smart kid Time is linear Memory is a stranger History is for fools Man is a tool in the hands Of the great God Almighty And they gave him command Of a nuclear submarine And sent him back in search of the Garden of Eden

Perfect Sense, Part II

Can’t you see It all makes perfect sense Expressed in dollars and cents Pounds shillings and pence Can’t you see It all makes perfect sense Little black soul departs in perfect focus Prime time fodder for the News at Nine Darling is the child warm in the bed tonight Hi everybody I’m Marv Albert And welcome to our telecast Coming to you live from Memorial Stadium It’s a beautiful day And today we expect a sensational matchup But first our global anthem Can’t you see It all makes perfect sense Expressed in dollars and cents Pounds shillings and pence Can’t you see It all makes perfect sense And here come the players As I speak to you now the captain Has his cross hairs zeroed in on the oil rig It looks to me like he’s going to attack By the way did you know that a submarine Captain earns 200,000 dollars a year Oh that’s less tax Marv yeah less tax Uh thank you Emery you’re welcome Now back to the game he fires one yes There goes two both fish are running The rig is going into a prevent defense Will they make it I don’t think so Can’t you see It all makes perfect sense Expressed in dollars and cents Pounds shillings and pence Can’t you see It all makes perfect sense Can’t you see It all makes perfect sense

The Bravery Of Being Out Of Range

You have a natural tendency To squeeze off a shot You’re good fun at parties You wear the right masks You’re old but you still Like a laugh in the locker room You can’t abide change You’re at home on the range You opened the suitcase Behind the old workings To show off the magnum You deafened the canyon A comfort a friend Only upstaged in the end By the Uzi machine gun Does the recoil remind you Remind you of sex Old man what the hell you gonna kill next Old timer who you gonna kill next I looked over Jordan and what did I see Saw a U.S. Marine in a pile of debris I swam in your pools And lay under your palm trees I looked in the eyes of the Indian Who lay on the Federal Building steps And through the range finder over the hill I saw the front line boys popping their pills Sick of the mess they find On their desert stage And the bravery of being out of range Yeah the question is vexed Old man what the hell you gonna kill next Old timer who you gonna kill next Hey bartender over here Two more shots And two more beers Sir turn up the TV sound The war has started on the ground Just love those laser guided bombs They’re really great For righting wrongs You hit the target And win the game From bars 3,000 miles away 3,000 miles away We play the game With the bravery of being out of range We zap and maim With the bravery of being out of range We strafe the train With the bravery of being out of range We gain terrain With the bravery of being out of range We play the game With the bravery of being out of range

Late Home Tonight, Part I

Standing at the window A farmer’s wife in Oxfordshire Glances at the clock it’s nearly time for tea She doesn’t see The phantom in the hedgerow dip its wing Doesn’t hear the engine sing But in the cockpit’s techno glow Behind the Ray Ban shine The kid from Cleveland In the comfort of routine Scans his dials and smiles Secure in the beauty of military life There is no right no wrong Only tin cans and cordite and white cliffs And blue skies and flight flight flight The beauty of military life No questions only orders and flight only flight What a beautiful sight in his wild blue dream The eternal child leafs through his War magazines And his kind Uncle Sam feeds ten trillion in Change into the total entertainment Combat video game And up here in the stands The fans are goin’ wild The cheerleaders flip When you wiggle your hip And we all like the bit when you take The jeans from the refrigerator and Then the bad guy gets hit And were you struck by the satisfying Way the swimsuit sticks to her skin Like BB gun days When knives pierce autumn leaves But that’s okay see the children bleed It’ll look great on the TV And in Tripoli another ordinary wife Stares at the dripping her old man hadn’t time to fix Too busy mixing politics and rhythm In the street below

[Various Arabic spoken (translation):

(Wife:) ”...And then what! Why don’t you ever help me? You all the time leave me alone at home and go join your fat friends in your endless useless discussions...I work for you and your family from dawn to dusk, and you don’t give a damn! I badly need to rest, I just wish the devil’s angel will soon come and take me with him...”

(Demonstrators:) ”...is Great;...is Great; God is great...Death, Death, Death to the imperialists...”

(TV commentator:) ”His days are rarely spent at home...as for her, she stays alone, she stays alone at home...while all the men are out gathered at the square, she’s left to loneliness and oblivion...” (explosion terminates the transmission)]

Late Home Tonight, Part II

Hark the wire service sing Clear the satellite link Check the fax machine Hold the lead story boys There’s some great pictures coming in Now the pilot’s heartbeat slows Palms dry out No questions only orders And the F-1 glides in nose-up Through the cloudbase and the Ground crew cheers as he puts down His landing gear Hey boy you’re a hero take this cigar And back home in Cleveland All the papers and the local TV stations Will be calling your ma And the farmer’s wife Shoos the cat off the chair She says sit down my dear Was the milking all right Our American friends are late home tonight

Too Much Rope

When the sleigh is heavy And the timber wolves are getting bold You look at you companions And test the water of their friendship with your toe And they significantly edge Closer to the gold Each man has his price Bob And yours was pretty low Now history is short the sun is just a minor star The poor man sells his kidneys In some colonial bazaar Que sera sera Is that your new Ferrari car Nice but I think I’ll wait for the F50 You don’t have to be a Jew To disapprove of murder Tears burn my eyes Moslem or Christian Mullah or Pope Preacher or poet who was it wrote Give any one species too much rope And they’ll fuck it up And last night on TV A Vietnam vet Takes his beard and his pain And his alienation twenty years Back to Asia again Sees the monsters they made In formaldehyde floating ’round Meets a gook on a bike A good little tyke [A nice enough guy] With the same soldier’s eyes What does it mean This tearjerking scene Beamed into my home That it moves me so much Why all the fuss It’s only two humans being It’s only two humans being Tears burn in my eyes What does it mean This tender TV This tearjerking scene Beamed into my home And you don’t have to be a Jew To disapprove of murder Tears burn my eyes Moslem or Christian Mullah or Pope Preacher or poet who was it wrote Give any one species too much rope And they’ll fuck it up

What God Wants, Part II

Do you believe in a better day Do you have faith in a golden way If you do then we must come together this day Come together as one united television audience Brought together by the sound of my voice United united financially united socially United spiritually and all possible ways Through the power of money And the power of your prayers What God wants God gets God help us all God wants dollars God wants cents God wants pounds shillings and pence God wants guilders God wants kroner God wants Swiss francs And God wants French francs Oui il veut des francs francais Ha ha God wants escudos God wants pesetas Don’t send lira God don’t want small potatoes God wants small towns God wants pain God wants clean up rock campaigns God wants widows God wants solutions God wants TV God wants contributions What God wants God gets God help us all What God wants God gets God wants silver God wants gold God wants his secret Never to be told God wants gigolos God wants giraffes God wants politics God wants a good laugh What God wants God gets God help us all God wants friendship God wants fame God wants credit God wants blame God wants poverty God wants wealth God wants insurance God wants to cover himself What God wants God gets God help us all

What God Wants, Part III

Don’t be afraid it’s only business The alien prophet sighed The vulture and the magpie took The cash box from its hook The monkey in the corner wrote The figures in his book Crazed the checkout lady’s fingers Flash across the till And the captain posts The menu of the day And in banks across the world Christian Moslem Hindu Jew And of every other race creed colour tint or hue Get down on their knees and pray The raccoon and the groundhog Neatly make up bags of change But the monkey in the corner Well he’s slowly drifting out of range Christ it’s freezing inside The veteran cries The hyenas break cover And stream through the meadow And the fog rolls in Through his bottle of gin So he picks up a stone That looks like a bone And the bullets fly And the rivers run dry And the fat girls sigh And the network anchor persons lie And the soldier’s alone In the video zone But the monkey’s not watching He’s slipped out to the kitchen To pile the dishes And answer the phone

Watching TV [Wa,Don Henley]

We were watching TV Watching TV We were watching TV Watching TV In Tiananmen Square Lost my baby there My yellow rose In her bloodstained clothes She was a short order pastry chef In a Dim Sum dive on the Yangtze tideway She had shiny hair She was the daughter of an engineer Won’t you shed a tear For my yellow rose My yellow rose In her bloodstained clothes She had perfect breasts She had high hopes She had almond eyes She had yellow thighs She was a student of philosophy Won’t you grieve with me For my yellow rose Shed a tear For her bloodstained clothes She had shiny hair She had perfect breasts She had high hopes She had almond eyes She had yellow thighs She was the daughter of an engineer So get out your pistols Get out your stones Get out your knives Cut them to the bone They are the lackeys of the grocer’s machine They built the dark satanic mills That manufacture hell on earth They bought the front row seats on Calvary They are irrelevant to me And I grieve for my sister People of China Do not forget do not forget The children who died for you Long live the Republic Did we do anything after this I’ve a feeling we did We were watching TV Watching TV We were watching TV Watching TV She wore a white bandanna that said freedom now She thought the Great Wall of China Would come tumbling down She was a student Her father was an engineer Won’t you shed a tear For my yellow rose My yellow rose In her bloodstained clothes Her grandpa fought old Chiang Kai-shek That no-good low-down dirty rat Who used to order his troops To fire on the women and children Imagine that imagine that And in the spring of ’48 Mao Tse-tung got quite irate And he kicked that old dictator Chiang Out of the state of China Chiang Kai-shek came down in Formosa And they armed the island of Quemoy And the shells were flying across the China Sea And they turned Formosa into a shoe factory Called Taiwan And she is different from Cro-Magnon man She’s different from Anne Boleyn She is different from the Rosenbergs And from the unknown Jew She is different from the unknown Nicaraguan Half superstar half victim She’s a victor star conceptually new And she is different from the Dodo And from the Kankabono She is different from the Aztec And from the Cherokee She’s everybody’s sister She’s symbolic of our failure She’s the one in fifty million Who can help us to be free Because she died on TV And I grieve for my sister

Three Wishes

Reached back for the bottle And rubbed against the lamp Genie came out smiling Like some Eastern tramp He said hey boy what’s happening What is going on You can have three wishes If you don’t take too long If you don’t take too long I said well I wish they were all happy in the Lebanon Wish somebody’d help me write this song I wish when I was young my old man had not been gone Genie said consider it done There’s something in the air And you don’t know what it is You see someone through the window Who you’ve just learned to miss And the road leads on to glory but You’ve used up your last wish Your last wish And you want her to come home Bring her home Genie said I’m sorry But that’s the way it goes Where the hell’s the lamp sucker It’s time for me to go Bye There’s something in the air And you don’t know what it is You see someone through the window Who you’ve just learned to miss And the road leads on to glory but You’ve used up your last wish Your last wish And you want her to come home

It’s A Miracle

Miraculous you call it babe You ain’t seen nothing yet They’ve got Pepsi in the Andes They’ve got McDonalds in Tibet Yosemite’s been turned into a golf course for the Japs And the Dead Sea is alive with rap Between the Tigris and Euphrates there’s a leisure centre now They’ve got all kinds of sports They’ve got Bermuda shorts They had sex in Pennsylvania A Brazilian grew a tree And a doctor in Manhattan saved a dying man for free It’s a miracle It’s a miracle It’s a miracle Another miracle By the grace of God Almighty And the pressures of the marketplace The human race has civilized itself It’s a miracle We got warehouses of butter We got oceans of wine We got famine when we need it We got designer crime We got Mercedes We got Porsche Ferrari and Rolls Royce We got choice She said meet me In the Garden of Gethsemene my dear The Lord said Peter I can see your house from here An honest man Finally reaped what he had sown And a farmer in Ohio has just repaid a loan It’s a miracle It’s a miracle It’s a miracle Another miracle By the grace of God Almighty And the pressures of the marketplace The human race has civilized itself It’s a miracle We cower in our shelters With our hands over our ears Lloyd-Webber’s awful stuff Runs for years and years and years An earthquake hits the theatre But the operetta lingers Then the piano lid comes down And breaks his fucking fingers It’s a miracle

Amused To Death

Doctor Doctor what is wrong with me This supermarket life is getting long What is the heart life of a colour TV What is the shelf life of a teenage queen Ooh western woman Ooh western girl News hound sniffs the air When Jessica Hahn goes down He latches on to that symbol of detachment Attracted by the peeling away of feeling The celebrity of the abused shell the belle Ooh western woman Ooh western girl The children of Melrose Strut their stuff Is absolute zero cold enough And out in the valley warm and clean The little ones sit by their TV screens No thoughts to think No tears to cry All sucked dry Down to the very last breath Bartender what is wrong with me Why am I so out of breath The captain said excuse me ma’am This species has amused itself to death Amused itself to death It has amused itself to death We watched the tragedy unfold We did as we were told We bought and sold It was the greatest show on earth But then it was over We oohed and aahed We drove our racing cars We ate our last few jars of caviar And somewhere out there in the stars A keen-eyed look-out Spied a flickering light Our last hurrah And when they found our shadows Grouped around the TV sets They ran down every lead They repeated every test They checked out all the data on their lists And then the alien anthropologists Admitted they were still perplexed But on eliminating every other reason For our sad demise They logged the explanation left This species has amused itself to death No tears to cry no feelings left This species has amused itself to death Amused itself to death

[(switch channels) (Alf Razzell:) ”Years later, I saw Bill Hubbard’s name on the memorial to the missing at Aras[?]. And I...when I saw his name I was absolutely transfixed; it was as though he was now a human being instead of some sort of nightmarish memory of how I had to leave him, all those years ago. And I felt relieved, and ever since then I’ve felt happier about it, because always before, whenever I thought of him, I said to myself, ’Was there something else that I could have done?’ (background: ”Put me down, put me down, I’d rather die, I’d rather die, put me down”) And that always sort of worried me. But having seen him, and his name in the register—as you know in the memorials there’s a little safe, there’s a register in there with every name—and seeing his name and his name on the memorial; it sort of lightened my...heart, if you like.” (woman:) ”When was it that you saw his name on the memorial?” ”Ah, when I was eighty-seven, that would be the year, ninety...eighty-four, nineteen eighty-four.”]

Checked against audio and lyrics sheet (SONY SBM CD edition) 950730

David Gilmour (25/5 1978)

Mihalis [instr](Gi)

There’s No Way Out Of Here [Gi](Gi)

There’s no way out of here When you come in You’re in for good There was no promise made The part you played The chance you took

There are no boundaries set The time and yet You waste it still So it slips through your hands Like grains of sand You watch it go

There’s no time to be lost You’ll pay the cost So get it right There’s no way out of here When you come in You’re in for good

There never was there an answer There an answer Not without listening Without seeing

There are no answers here When you look out You don’t see in There was no promise made The part you played The chance you took

There’s no way out of here When you come in You’re in for good

There never was there an answer There an answer Not without listening Without seeing

There’s no way out of here When you come in You’re in for good There was no promise made The part you played The chance you took

There’s no way out of here When you come in You’re in for good There’s no way out of here When you come in You’re in for good There are no answers here When you look out You don’t see in There’s no way out of here When you come in You’re in for good

Cry From The Street [Gi](Gi,E. Stuart)

Black shadows dance Beyond the neon glow There’s rumors in your head And just one way to go Oh so tight It makes your blood run cold The way is blind Try to save your soul Don’t even pray Just cry from the street

It’s fine as you leave The one they keep at home So you love someone That you thought you owed You know it’s a crime But what can you do? That’s the price you pay With each turn of the screw But don’t pay no mind Just cry from the street

You really love the night Even though you’re alone It’s such a cryin’ shame That things are so wrong It’s the sign of the scream That it blinds your feet I guess the chance is ours You wish to stay and sleep But don’t even pray Just cry from the street

So Far Away [Gi](Gi)

What am I suspended here? Should I kid myself that I have no fear? I get no choice, I just have to wait It may already be too late I don’t know what’s going on She is here, but I am gone I am gone

My love is lying here She’s far away though she looks so near ”Time will tell” she says to me When a warm coat is all I need Her peace of mind, her strentgh of will It will come she’s sure it will And how can I put my mind at rest? I feel I’m coming up second best

And I’m caught with my own life Is this a dream, or is it real life? Oooh sleep, come on me soon I can’t bear this lonely room

Look at that hand lying there The room is dark she shows no fear I’m lying still, my eyes are wide My heart is pumping, I’m still alive I’m still awake against my will What will it ever take To still this burning in me?

And I’m caught with my own life Is this a dream, or is it real life? Ooooh, sleep, come on me soon I can’t bear this lonely room Sleep, come on me soon I can’t bear this lonely room Sleep, don’t take too long Can’t keep singing this lonely song

Short And Sweet [Gi](Gi,R. Harper)

You ask What is the quality Of life? Seeking to justify the part you play And hide Fearing it incomplete To try To make it any more or less than short and sweet

But short Short is from you to me As close As we are wont to try to make it be We’re caught Watching the dark in the sky Who knows? Helpless it’s time and self to hold the time of day

And you You are a fantasy A view From where you’d like to think the world should see Be true And you will likely find A few Building a vision new and justice to our time

And we We the immoral men We dare Naked and fearless in the elements And free Carefree of tempting fate Aware And holding off the moral nightmare at the gates

And sweet Sweet as a mountain stream We’ll look Toward a new day breaking in the east And meet As every future dream Unfolds And surely quality that is the very least

Raise My Rent [instr](Gi) No Way [Gi](Gi)

There’s no way I’m gonna let go There’s no way Because it’s my show I’m hangin’ on For a little while I won’t go down easy That’s not my style

I’m all tied up Tied up in a knot And I can’t decide Just what it is I’ve got Did I get out of touch Did I lose my way I’ve not forgot No, not a single day

Now you might believe I’m not a happy man Putting myself apart From the common clan What a boat we’re sailing We had a week or two But I know it sounds Like me and you

So there’s no way I’m gonna let go There’s no way Because it’s my show I’m hangin’ on For a little while I won’t go down easy That’s not my style

It’s Deafinitely [instr](Gi) I Can’t Breathe Anymore [Gi](Gi)

I can’t breathe anymore Why that is I’m not sure I’ve got my feet on the ground In fact I’m flat on the floor My direction is lost I should pause to take stock It’s a new one on me To even dream that you’re free

It’s a tale without an end But the end is there for all to see And if the end weren’t so real Coming around and ’round again on me I guess it’s all about my pride I want to be there if it kills But with or without God on my side I know that I really will

About Face (1984)

Until we Sleep [Gi](Gi)

Dance, to the movement of the stars Sing till the walls around us ring Pray that it never fades away Until we sleep

Feast, go ahead now, take your fill Drink, till the room begins to sway Play, and your cares will blow away Until we sleep Into a deep and dreamless sleep our spirits fall And what we sow we cannot reap nor keep at all

Work, though the work will break your back Rest, hope to feel your spirits rise But the years, have cut you down to size Now go to sleep Into a deep and dreamless sleep our spirits fall And what we sow we cannot reap nor keep at all

Murder [Gi](Gi)

Some of them standing, some were waiting in the line As if there was something that they thought they might find Taking some strength from the feelings that always were shared And in the background, the eyes that just stared

What was it brought you out here in the dark? Was it your only way of making your mark Did you get rid of all the voices in your head? Do you now miss them and the things that they said?

On your own admission you raised up the knife And you brought it down ending another man’s life When it was done you just threw down the blade While the red blood spread wider like the anger you made

I don’t want this anger that’s burning in me It’s something from which it’s so hard to be free But none of the tears that we cry in sorrow or rage Can make any difference, or turn back the page

Love on the Air [Gi](Gi)

Love on the air I keep transmitting but reception is hazy I don’t get an answer Keep sending it faster Always knew it was crazy To put my love on the air.

No one will hurt me again No one will cause me to lie No one will control me by pain No one will cause me to cry I was looking for love In wandering eyes Like a ship trying to fix on a beacon I learned how to sigh On the ribbon and wires It’s a habit that’s so hard to weaken

No one will ever manipulate Make me promise to do or die No one can make me hesitate What can I lose if they try I was looking for love Like the very first time Didn’t realise love never left me Contradicted - conspired I connived and designed Nothing on earth could arrest me.

Reception is hazy When you put your love on the air Always knew it was crazy To put my love on the air But I only communicate When I put my love on the air You don’t have to consummate Love on the air

Blue Light [Gi](Gi)

She’s a blue light, yes it cannot be denied She will give you back your pride She squeezes, she freezes, she does what she pleases But she says she thinks of others, all the time

You must rise to her defense when she’s in danger Turn around and you’re looking at a stranger She steals your savings from under your bed And leads you into places even angels fear to tread She’s a blue light

She’s not amused, when you’re trying to upstage her So you choose to do nothing to enrage her ’Cause under her mantle you feel safe from the cold And you’re safe from the danger of ever growing old She’s a blue light

[I’m not very good with this funky sort of thing, But, uh, sometimes, when the inspiration strikes, I can get down with the best of them.]

Out of the Blue [Gi](Gi)

Out of the blue on the wings of a dove A messenger comes, with the beating of drums It’s not a message of love

Our childern are born, and we keep them They must have the right, to live in the light To be safe from the storm

Out of the blue, with wings on his heels A messenger comes, bearing regrets For the time that he steals

But steal it he will, my children’s and mine Against our desires, against all our needs Our blood spilled like wine

Over and over we call … no one hears And further and further and further we fall And though we pray that we soon will awake It is clear, that it’s no dream at all Our lives are at stake

I cannot believe, nor even pretend That the thunder I hear, will just disappear And the nightmare will end

So hold back the fire, because this music is true When all’s said and done, the ending will come From out of the blue

All Lovers are Deranged [Gi](Gi,Townshend)

It takes a fool to phone a fool When both have said it all We make the rule, bemoan the rule That neither one should call But love that was Is love that is Demands to always be unchanged But then all lovers are deranged.

We walk away with memories And clutch them to our hearts We’re disembodied entities We move in fits and starts For burning wine Intoxicates, And takes all caution in its flames All lovers are deranged.

You know that you don’t really fall in love Unles you’re seventeen The break of day will make your spirits fly But you can’t know what it means Unless you’re seventeen.

It takes a fight to start a fight And differences remain We have the right to think we’re right We’re addicts feigning shame For love recalled Is love reborn We’re determined to relive the pain But then lovers are deranged.

You Know I’m Right [Gi](Gi)

You can scream and shout with all your might Dig in your heels and hold on tight Either you are wrong or I am right

You speak the lines you’ve overheard The ring of truth in every word You know you’re right and that’s absurd

We really seem to have a problem here But it is you or me Whatever I have going through my mind You always have to disagree It’s just a matter of opinions It’s not a simple fact Why don’t you try to see the other side Don’t turn your back

Now we survey this silent battleground Recriminations all around And still no compromise is found

Now we really have a problem And it won’t just disappear And all the friends we thought we could rely on Just want to whisper in my ear

”It’s just a matter of opinions You know you keep both in sight Why should you bother with the other side When you know yours is right.”

Cruise [Gi](Gi)

Cruise you are making me sing Now you have taken me under your wing Cruise, we both know you’re the best How can they say you’re like all the rest

Cruise, we’re both travelling so far Burning out fast like a shooting star Cruise I feel sure that your song will be sung And will ring in the ears of everyone

Saving our children, saving our land Protecting us from things we can’t understand Power and Glory, Justice and Right I’m sure that you’ll help us to see the light And the love that you radiate will keep us warm And help us to weather the storm

Cruise, you have taken me in And just when I’ve got you under my skin You start ignoring the fears I have felt ’Cause you know you can always make my poor heart melt

Please don’t take what I’m saying amiss Or misunderstand at a time such as this Because if such close friends should ever fall out What would there be left worth fighting about

Power and glory, justice and right I’m sure that you’ll help them to see the light Will you save our children, will you save our land And protect us from all the things we can’t understand? Power and glory and justice for all Who will we turn to when your hard rain falls?

Let’s get Metaphysical [instr](Gi)

Near the End [Gi](Gi)

And when you fell you’re near the end Will you just turn it over and start again Is there a stirring in your heart As the time comes when we will have to part?

And when you fell you’re near the end And there’s a stranger where once was a friend And you are left without a word Only the whispers that you’ve overheard

Standing in silence, holding my breath Disconnected and dry And though I’m certain that there’s nothing left To hold on to, to give or to try Some things never change, no don’t ever change And I’m feeling the cold Thinking that we’re getting older and wiser When we’re just getting old

And when you feel you’re near the end And what once burned so bright is growing dim? And when you see what’s been acheived Is there a feeling that you’ve been deceived?